### v1

# Dialects for CoAP-like Messaging Protocols

Abstract
Messaging protocols for resource limited systems such as distributed IoT systems are often vulnerable to attacks due to security choices made to conserve resources such as time, memory, or bandwidth. For example, use of secure layers such as DTLS are resource expensive and can sometimes cause service disruption. Protocol dialects are intended as a light weight, modular mechanism to provide selected security guarantees, such as authentication. In this report we study the CoAP messaging protocol and define two attack models formalizing different vulnerabilities. We propose a generic dialect for CoAP messaging. The CoAP protocol, dialect, and attack models are formalized in the rewriting logic system Maude. A number of case studies are reported illustrating vulnerabilities and effects of applying the dialect. We also prove (stuttering) bisimulations between CoAP messaging applications and dialected versions, thus ensuring that dialecting preserves LTL properties (without Next) of CoAP applications.
1 Introduction
There is a rapidly growing number of networks of IoT devices that impact our daily lives. They enable smart homes, offices, factories, and infrastructure. They are increasingly used in health care, precision agriculture, logistics, supply chain management, and situation awareness. The IoT devices sense and act on the physical environment and must operate using limited resources (energy, bandwidth, memory, compute power, …). The vulnerabilities of small inexpensive devices are magnified when deployed at scale. Security is crucial for safe operation but security is typically resource intensive. Design of such systems must balance balance resources used for messaging and for security.
Protocol dialects are a light weight mechanism that can provide authentication and possibly additional security services such as integrity. A dialect transforms the underlying protocol to mutate messages before sending such that only dialect partners can revert the mutation for delivery to the intended receiver. A key feature of dialects is to rapidly change the mutation used to prevent attackers from using any decoding information they might gain. This moving target defense means that the complexity of mutation processing can be kept low without compromising the security guarantee. The need to change mutations presents a significant challenge to synchronize change across distributed network nodes.
Existing dialect works, for example [sjoholmsierchio-etal-2021netsoft, gogineni-etal-2022arXiv, ren-etal-2023seccom], usually consider protocols running over TCP or other reliable transport. In this case mutations can reasonably be synchronized based on time. In the case of unreliable transport where messages can be delayed, reordered, lost, or repeated, time-based synchronization is problematic. Unreliable transport is mentioned in [GEMM-2023esorics] as one classification parameter, but synchronization mechanisms for this case are not treated.
In this work we study dialects for protocols running over unreliable transport, using the CoAP messaging protocol [RFC7252] as prototypical example. We investigate the use of counters, analogous to those used to prevent replay in security protocols. for synchronizing. We consider two methods of evaluating the CoAP specification and dialect wrapper. One is based on the CoAP vulnerabilities discussed in [coap-attacks] and the other is based on properties a CoAP enabled endpoint could assume when running over an unreliable transport layer such as UDP.
For the vulnerability case, we formalize the attack scenarios presented in [coap-attacks] as motivation for strengthening the specification. These scenarios consider an attacker that can only block/drop or delay messages. We use reachability analysis to look for attacks. In some cases the Maude endpoints exhibit the expected vulnerability. in some cases because the Maude specification implements some of the stronger mechanisms, the attack fails, but there are related attacks. In some cases an attacker that can redirect messages can realize the attack, but dialecting mitigates such attacks.
For the CoAP properties case the goal is to identify guarantees provided by CoAP messaging running on an unreliable transport, identify attack points, and corresponding attack model, and prove that dialecting provides adequate defense. This is formalized using stuttering bisimulation relations.
Contributions.
The main contributions of this report are:
  • Executable specification of the CoAP messaging protocol (Section 3).
  • Specification of two families of resource limited attack models (Section 4): an active attacker that can modify messages in transit and a reactive attacker that can observe messages in transit and transmit (modified) copies.
  • Specification of a generic dialect layer and transform for messaging protocols (Section 7).
  • Reachability analysis demonstrating CoAP vulnerabilities identified in [coap-attacks] (Appendix 0.B). These require an active attacker.
  • Reachability analysis demonstrating diverse reactive attacks (Section 5).
  • An analysis of protection provided by the proposed dialect class in unreliable/reliable networks (Section 8).
  • Proof of stuttering bisimilarity theorems characterizing effects of the dialect transformation (Appendix 0.C).
Plan.
Section 2 provides background information on CoAP, dialects, and rewriting logic/Maude. Section 3 describes the Maude specification of CoAP messaging. Resource limited attack models are presented in Section 4 with examples illustrating a variety of attacks presented in Section 5. In Section 6 we present the abstract mutation functions of the proposed dialect scheme. The Maude specification of the CoAP dialect is given in Section 7. In Section 8 we discuss the properties of dialects such as the CoAP dialect. The relation to [GEMM-2023esorics] is discussed in Section 9. Concluding remarks are given in Section 10.
In Appendix 0.A we spell out key properties of CoAP system executions, including characterizing the minimal time elapse function. In Appendix 0.B we specify the scenarios illustrating CoAP vulnerabilities listed in [coap-attacks] and do reachability analysis to determine which vulnerabilities are realized by the Maude specification. We also analyze the dialected versions of each scenario. In Appendix 0.C we prove two bisimulation theorems for dialected CoAP systems, verifying relations proposed in Section 8.
2 Background
2.1 The CoAP protocol
The Constrained Application Protocol (CoAP) [RFC7252, coap-attacks] is an HTTP-like client-server Protocol for use by resource-constrained devices (e.g. low power) and networks (low bandwidth, lossy). CoAP provides a request/response RESTFUL interaction model between application endpoints. Servers hold resources that can be generated, updated, accessed by clients. An endpoint can be a client, a server, or both. A design goal is to keep message overhead small, to limit need for fragmentation. Note that server resources may be connected to actuators having effects on the environment.
CoAP features include UDP binding (reliability supported at the protocol level by a retry mechanism) and asynchronous two-way message exchange. A binding to Datagram Transport Layer Security (DTLS) provides security, but with substantial overhead. Expected applications include: smart buildings, instrumented infrastructure, medical devices, etc.
RFC7252 specifies a binary format for CoAP messages. Here we present an abstract data-type view. A CoAP message consists of four parts: Header, Token, Options, and Payload. The header has four parts: Version, Type, Code, and MsgId (we will ignore the version).
The header type is one of {CON, NON, ACK, RST}. A message of type CON is confirmable. The receiver must acknowledge with a message of type ACK, in addition to a response if appropriate. The sender should resend a CON message if no ACK is received within ACK-TIMEOUT time. There is a bound on the number of resends (a configuration parameter). A message of type NON is non-confirmable. If received it may require that a response is sent, but no separate acknowledgment is sent. A message of type RST is a reset message, used to indicate receipt of message that can’t be handled, also use as a kind of ping.
Code is an HTTP like code that determines if the message is a request or response. In the case of a request the code also determines the method–one of GET,POST,PUT,DELETE. In the case of a response, the code determines if the result is a success (with value in the payload), a client error, or a server error.
The MsgId header element is a string that uniquely identifies the message (among messages from the sender). It is used to match acknowledgements to CON messages. It also is used to provide limited replay protection. Message identifiers have a limited lifetime (a configuration parameter) after which they can be reused.
The token message component is a unique identifier generated by a requestor to match request to response. Options is a list of name-value pairs used to specify additional information, including: content-format, request Uri, and response Uri (of a created resource). Payload contains additional request information (for example the value to PUT) or response value (for example the result of a GET).
The rules for handling requests with a given method are similar to those for HTTP. A response can be piggy backed on the acknowledgement of a CON request. A separate response can be of type CON or NON. Servers are required to respond to legitimate requests, but they should only process a request once. Thus if a confirmable request is resent, for example because the acknowledgement was lost or delayed, the server may just resend the ACK or it may resend the actual response, but it should not repeat the actions needed to carry out the response.
2.2 Dialects
A protocol dialect is a variation on protocol messages intended as a light weight mechanism to provide some additional security properties such as forms of authentication or message integrity.
A dialect uses a family of lingos to transform protocol messages. A lingo provides a pair of functions that are inverse of one another, possibly modulo parameters. Dialects provide moving target defense in the sense that lingo parameters and choice of lingo change over time in a dialect specified manner. This means weaker forms of message transformation can be used, because even if a message transformation is determined, it doesn’t help as that transformation will not be used on later messages. It also means that honest participants must synchronize on the choice of lingo and lingo parameters.
A typical application of dialects is to networks where at least some nodes are resource limited devices. An example would be smart infrastructure with sensors and actuators running on battery power, and with limited memory and cpu capability. Thus a dialect should not require substantial resources. In addition to providing new security properties, a dialect should preserve important properties of the underlying protocol. The protocol may be running on a network with limited bandwith, thus a dialect should not generate significant additional traffic, and should not significantly increase message size.
The following are some examples of dialects and their application. Dialects to improve SDN (Software defined network) security are defined and evaluated in [sjoholmsierchio-2019nps, sjoholmsierchio-etal-2021netsoft]. Two transformations based on HMAC over the message data are proposed. One places the HMAC in the session id slot. This limits the HMAC size and thus the hash key is changed every second. For messages that use this slot, a 512-bit HMAC tag is appended to data packets. Sequence numbers are included in the hashed content to prevent replay. Keys are generated from pre-shared symmetric keys, delivered out-of-band. Dialects are implemented using proxies inserted into the communication path to carry out the message transformations. One goal is to protect against TLS downgrade attacks by providing authentication for initial TLS communications.
Dialects for FTP and MQTT protocols using bit shuffling or packet splitting are studied in [gogineni-etal-2022arXiv]. A self-synchronization mechanisms is propsed. In [gogineni-etal-2021site] the authors propose a Deep Neural Network mechanism for synchronization. A dialect is implemented by (conceptually) inserting a separate process between the protocol and the network.
Dialecting is proposed in [ren-etal-2023seccom] to mitigate known exploitable vulnerabilities in IoT devices deployed in large numbers thus enabling massive attacks. The idea is to transform functions that construct and parse message data structures to construct and parse mutated data structures. The message constructor/parser functions are identified automatically and message mutations are typically per message field. Specific mutations are designed using testing. Dialects are implemented by modifying the identified message constructor and parser functions of the protocol implementation. The method is evaluated on implementations of MQTT, LwM2M, and DDS. In the experiments, the mutations successfully blocked reproducible attacks corresponding to a selected set of CVEs reported for the protocols
The above works propose dialects and evaluate them experimentally. In [GEMM-2023esorics] a formal framework for specifying and applying dialects is proposed. The ideas are illustrated using the MQTT protocol. This work and its relation to the current work is discussed in detail in section 9.
2.3 Rewriting logic and Maude
[…]
Maude [clavel-etal-07maudebook] is a language and tool implementing rewriting logic. Maude can be downloaded from the Maude website [maudeweb]. Rewrite theories are represented by Maude modules. Functional modules (declared using the keyword fmod at the beginning and endfm at the end) represent equational theories (with no rules). System modules (declared using the keyword mod at the beginning and endm at the end) represent rewrite theories with rules. We introduce the main Maude syntax elements with examples from the CoAP specification (discussed in detail in Section 3). Sorts and their ordering are declared using the keywords sort and subsort. For example, the following declares two sorts Msg (the message sort) and MsgS (multisets of messages) with Msg (singleton sets) being a subsort of MsgS.
    sort Msg MsgS .
    subsort Msg < MsgS .
Functions/operators are declared using the keyword op. For example, the operator m constructs a message from two strings (the target and source) and a term of sort Content. The attribute ctor says that m is a constructor and will not be subject any equations. mtM is a constant (no arguments) denoting the empty message set. The operator __ has empty syntax, the blanks denoting argument positions. It correspond to union of multi sets. That union is associative and commutative with unit the empty set is expressed by the attributes assoc, comm, and id: mtM. The ctor attribute says that no other equations apply to elements of MsgS. Thus multisets are generated from the empty set and singletons by union.
    op m : String String Content -> Msg [ctor] .
    op mtM : -> MsgS [ctor] .
    op __ : MsgS MsgS -> MsgS [ctor assoc comm id: mtM] .
This is a general pattern used to specify (multi)sets of a given sort. A similar pattern is used to specify lists of elements, omitting the comm attribute.
Equations are introduced with the key word eq (or ceq for conditional equations). For example, the function getTgt, that gets the target of a message, is defined by the following declaration and equation.
    op getTgt : Msg -> String .
    eq getTgt(m(tgt:String,src:String,c:Content)) = tgt:String .
Here we use inline variable declaration. For example, tgt:String declares tgt to be a variable of sort String. Alternatively, we could declare the variable globally.
    vars tgt src : String .
    var c : Content .
    eq getTgt(m(tgt,src,c)) = tgt .
Rewrite rules are signalled by the keyword rl (crl for conditional rules), optionally followed by a label in []s. The rule labeled net applies to any configuration that has a network element of the form net(dmsgs0 dmsg,dmsgs1), and moves dmsg from the first to the second argument of the constructor net.  [^1] A delayed message dmsg has the form msg @ d where d is a natural number representing the amount of time it will take the message to traverse the network.
    rl[net]:
    net(dmsgs0 dmsg,dmsgs1) => net(dmsgs0,dmsgs1 dmsg)
[…]
There are many Maude tools for static and dynamic analysis of a system specification. In the present work we only need the rewrite command for prototyping and the search command for reachability analysis. Assume you have defined a module SCENARIO that includes all the modules defining rules and data structures defining your model. Assume further that system configurations are of sort Sys and the module SCENARIO defines a constant initSys of sort Sys. Then we can use the command
    rew [n] initSys .
to find one way that initSys might evolve for n steps. If you omit [n] the execution will run until a state is reached that has no successors (a terminal state). This should be used with care as there may not be a reachable terminal state! If a boolean function P(sys:Sys) is defined, one can use the search command to carryout reachability analysis. The following search command will return the first n distinct reachable states that satisfy the property P.
  search [n] initSys =>+ sys:Sys such that P(sys:Sys) .
If [n] is omitted Maude will search for all reachable states satisfying P. Again caution is suggested as there may be infinitely many such states.
Example initial configurations and properties for our CoAP specification are defined in section 3 and reachability analysis experiments for different attack models are given in Sections 5 and Appendix 0.B.
3 Executable Specification of CoAP
The Maude specification of the CoAP messaging protocol consists of specifications of data structures to represent execution state 3.1, rules specifying the protocol behavior 3.2, and an attack specification framework 3.3. The data structures include a message data type and data structures representing endpoint and network state. The basic CoAP rules model sending and receiving messages, flow of messages in the network, and passing of time. The attack model consists of a specification of attacker capabilities and a rule for application of a capability to modify communications. The Maude specification of the two families of attack models is given in Section 4.4 The CoAP specification concludes with definitions of functions for generating application message lists and scenario initial states ( Section 3.4) and some small example scenarios and analyses to illustrate the ideas (Section 3.5).
3.1 Data types
Message data type.
The specification of the message data type is a direct formalization of the (abstract) structure of messages as specified in [RFC7252]. Messages (sort Msg) are constructed by the operation
    op m : String String Content -> Msg [ctor] .
thus a typical message term has the form m(tgt,src,content) with tgt, src being strings identifying the receiver (target) and sender (source) respectively.
The sort DMsg of delayed messages consists of terms of the form msg @ d where msg is of sort Msg and d is the delay, a natural number (sort Nat). The delay is used to model the time between sending and receiving a message (network latency). It is also used to model the timer controlling resend of a confirmable message.
As discussed in Section 2.1 message content (sort Content) is constructed from sorts Head, multisets of Options, and Body (aka payload).
    op c : Head String Options Body -> Content [ctor] .
We define a constant mtBody of sort Body to stand for the empty payload, and assume non empty payload is represented by strings: b(str). Elements of sort Head are constructed by
    op h : String String String -> Head  [ctor] .
where in h(type,code,mid) the first string is the message type ("CON", ‘‘NON’’, "ACK" or "RST"), the second string is the HTTP-like code, and the third string is the message unique identifier, generated by the sender, as discussed in 2.1
The String element of a message content is the token used to match a response to the corresponding request. Each request from a given source to a given target should have a unique token generated for it.
Sort Options is a multiset of sort Option where mtO is the empty option set. An individual option has the form o(oname,oval) where oname is a string naming the option and oval is the option value of sort String or Nat. The current specification supports options representing request URIs and URIs in a response resulting from creation of a resource. For example the equation
    eq getPath(opts o("Uri-Path",path)) = path .
defines how to access the Uri-path of a request. Multisets of delayed messages (sort DMsgS, id: mtDM) are defined similarly to multisets of messages shown above.
Selectors are defined for each component of a message. For example the selector for message type is defined as follows:
      op getType : Msg -> String .
      op getType : Content -> String .
      op getType : Head -> String .
      eq getType(m(dst:String,src:String,c:Content))
                = getType(c:Content) .
      eq getType(c(h:Head,tok:String,opts:Options,body:Body))
                = getType(h:Head) .
      eq getType(h(type,code,mid)) = type .
The CoAP protocol itself does not generate requests, rather it packages and transmits application level messages. For testing and analysis purposes we represent the application by a list of application level messages (requests) (sorts AMsg, AMsgL) to be transmitted. An application message consists of six strings and a message body. The strings specify: the application id, the message target and type, the method, and the resource path and query parameters if any.
   ****     appid    tgt   type   meth   path  qparams body
  op amsg : String String String String String String Body
          -> AMsg [ctor] .
The empty AMsgL is nilAM and list concatenation is _;_.
Endpoint state.
A CoAP network system consists of a set of endpoints (devices running CoAP) and the network. The network is modeled as a pair of delayed message sets net(dmsgs0,dmsgs1). Newly sent messages enter the first component, are moved by the net (or an attacker) to the second component from which they are delivered once the delay is zero.
An endpoint (also referred to as a device) is a term of sort Agent of the form
        [epid | attrs]
where epid is a string identifying the endpoint and attrs is a set of attributes representing the endpoints current state. CoAP endpoint attributes include attributes to track status of msgs sent and received in order to implement the rules for reliability (confirmable messages) and deduplication. They also include configuration parameters allowing each endpoint to be configured differently if desired.
  • w4Ack(dmsgs): dmsgs is the set of confirmable messages sent and not yet acknowledged. The delay part is the amount of time to wait before resending the message;
  • w4Rsp(msgs): a message in msgs is either a non-confirmable request waiting for a response, or a confirmable request for which an ACK has been received but the response is still pending;
  • rspRcd(msgs): logs responses received to avoid reprocessing repeated responses;
  • rspSntD(dmsgs): logs responses sent where the delay represents the message id/token lifetime. It can be used to resend the response in case the original is lost, and to detect replay attempts within the message id lifetime.
  • rsrcs(rmap): rmap is a map from resource paths to values, representing server resources state;
  • ctr(n): n is a natural number used to model generating fresh/unique message Id strings (genMid(prefix,n)) and unique tokens (genTok(prefix,n));
  • sendReqs(amsgl): a list of application messages to be transmitted
  • config(cbnds): a map from global parameter names to configured values, examples include message send delay (abstracts time in transit, cb("msgSD",n)), maximun number of resends (cb("MAX_RETRANSMIT", n0)), backoff parameter for computing delay between resends (cb("msgQD",n1)), congestion control parameters such as time to wait for an ACK before resending a confirmable message (cb("ACK_TIMEOUT", n2)), etc..;
  • toSend(dmsg): an attribute used to store information returned by auxiliary functions
A number of auxiliary functions on attributes are defined for use in specifying rules. We discuss two examples. The function getMsgSndDelay looks up the "msgSD" configuration parameter. This parameter models average network transit time of a message.
    op getMsgSndDelay : Attrs -> Nat .
    op getMsgSndDelay : CBnds -> Nat .
    eq getMsgSndDelay(attrs config(cbnds)) =  getMsgSndDelay(cbnds) .
    eq getMsgSndDelay(cb("msgSD",n) cbnds) = n .
The function findRspRcd checks if a response message from dst has been received in response to a request with identifier mid and token tok. It uses the message selectors getSrc which would be the destination of the corresponding request, and getTok to use to match the response to a request. findRspRcd is used to avoid reprocessing responses and also to be able to resend and ACK to a confirmable response.
    ****                    dst  mid     tok
    op findRspRcd : Attrs String String String -> Bool .
    op findRspRcd : MsgS String String String -> Bool .
    eq findRspRcd(devatts rspRcd(msgs), dst,mid,tok) =
          findRspRcd(msgs, dst,mid,tok) .
    ceq findRspRcd(msgs msg, dst,mid,tok) = true
      if getSrc(msg) == dst
      /\ getTok(msg) == tok .
    eq findRspRcd(msgs, dst,mid,tok) = false [owise] .
3.2 Rules
Five rules are used to specify CoAP executions. The rules labelled devsend and rcv specify transmission and receipt of messages by an endpoint. The rule labelled ackTimeout specifies when a confirmable message is resent. The rule labelled net specifies the network action of moving a delayed message from the input side to the output side. Finally, the rule labelled tick specifies the passing of time. Rules devsend, rcv and ackTimeout operate on sub-configurations consisting of an endpoint and the network, the rule net simply transforms the network configuration element, while tick rule requires the full system configuration {conf}, formed by encapsulating a configuration term in curly braces. The net rule was discussed in Section 2.3. The remaining rules are discussed in the remainder of this (sub)section.
Sending a message.
The rule devsend sends the first element, amsg, of the application message list attribute. The variable devatts matches the remaining attributes of the sending endpoint state. The rule can only fire if noW4Ack(devatts) is true and sending is enabled (canSend(devatts)). These tests model the CoAP congestion control. noW4Ack(devatts) holds if the number of confirmable messages awaiting acknowledgements is not greater than the configuration parameter "w4AckBd" and canSend checks that the sndCtr attribute of devatts is 0. The function sndAMsg constructs the delayed message to send, dmsgs, and updates devatts. To construct dmsgs, a message id and token are generated and the method string is converted to a code, then the header, options and body are used to produce the message content. The message delay is obtained from the configuration parameter "msgSD". If the message is type CON it is added to the w4Ack attribute of devatts with a delay given by the configuration parameter "ACK_TIMEOUT", otherwise the message is added to the w4Rsp attribute. The ctr attribute is incremented by 2 (the two uses to generate a message id and a token), and the sndCtr is reset using the configuration parameter "msgQD".
    crl[devsend]:
      [epid | sendReqs(amsg ; amsgl) devatts ]
      net(dmsgs0,dmsgs1)
      [epid | sendReqs(amsgl) devatts1]
      net(dmsgs0 dmsgs,dmsgs1)
    if noW4Ack(devatts)
    /\ canSend(devatts)
    /\ devatts1 toSend(dmsgs) := sndAMsg(epid,amsg, devatts) .
Receiving a message.
A message with 0 delay is received by the target endpoint using the rule labelled rcv. The work is done by the function rcvMsg which uses the message header code to classify the message as “Request”, “Response”, “Empty”, or “UnKnown” and calls rcvRequest, rcvResponse, rcvEmpty, or sendReset (if confirmable). Otherwise the message is dropped. An empty message is either an acknowledgement or a reset. In the acknowledgement case, the corresponding delayed message in the w4Ack attribute is removed if any, and the underlying message is added to the w4Rsp attribute to be able to process the pending response. In the reset case, if the message is confirmable an acknowledgement is sent (this implements a “PING” functionality) otherwise the reset is ignored.
    crl[rcv]:
      [epid | devatts] net(dmsgs0,dmsgs1 msg @ 0 )
      [epid | devatts1 ] net(dmsgs0 dmsgs,dmsgs1)
    if getTgt(msg) == epid
    /\ toSend(dmsgs) devatts1 := rcvMsg(epid, devatts, msg) .
The rules for handling requests are specified in [RFC7252] section 5.8. The function rcvRequest implements these rules. It first checks whether it has already sent a response to this request using matchingRsp(devatts,msg). If so, if the request is confirmable, the ACK is resent, otherwise the message is ignored. If the request is new the method is computed from the code and the appropriate method specific function is called to process the message. As an example, if the method is "GET", the function rcvGet is called. This function sends a single message in response:
    msg = m(src,epid,c(h(rtype,code,rmid),tok,mtO,body) .
rmid is a new message id, src is the sender of the request, tok is the token of the request message. The type, rtype, is "ACK" if the request in confirmable. In this case, msg is a combined acknowledgement and response. The type is "NON" if the request is nonconfirmable. In this case msg is a simple response. body is the result of attempting to access the resource at the path specified in the request message options using getResourceVal(devatts,path). It is b(val) if a value val is found, and mtBody if the resource does not exist. The message code is "2.05" if a resource value is found, indicating successful GET. The message code is "4.04" if the resource is not found, indicating an error. rcvGet updates the endpoint attribute by incrementing the counter, ctr, and recording that the request has been processed by adding msg to the rspSnt attribute.
The function rcvResponse checks if it is combined with an acknowledgement or a simple response. In the former case, the matching request is removed from the w4Ack attribute, if any, and adds the response message to the rspRcd if it has not already been seen. This abstracts interaction with the application layer. In the case of a simple response, if a response has already been received, an acknowledgment is sent if the response is confirmable and otherwise the message is ignored. If this is a new response, it is recorded in the rspRcd attribute, any occurrence of the matching request is removed from the w4Ack and w4Rsp attributes. If the response is confirmable an acknowledgement is sent.
ReSending a message.
The rule ackTimeout can fire if a message in the w2Ack attribute has delay 0. The option o("rcnt",n) records the number of resends of the message. If this count is less than the max allowed (configuration parameter "MAX_RETRANSMIT") then the message is returned to the w4Ack attribute with a new ack wait time, delay, computed using the backOff function that doubles the delay for each resend. Finally, the message with normal sending delay is put in the network input side.
  crl[ackTimeout]:
    [epid | w4Ack((msg @ 0) dmsgs) config(cbnds) attrs]
    net(dmsgs0,dmsgs1)
    [epid | w4Ack(dmsgs2 dmsgs) config(cbnds) attrs]
    net(dmsgs0 dmsg, dmsgs1)
  if  m(dst,epid,c(hd,tok,opts o("rcnt",n),body)) := msg
  /\ opts0 :=  o("rcnt",s n )
  /\ delay := backOff(getAckWait4(cbnds), s n)
  /\ dmsg :=  m(dst,epid,c(hd,tok,opts,body)) @
                getMsgSndDelay(cbnds)
  /\ dmsgs2 :=  (if n < getMaxReSnd(cbnds)
                 then m(dst,epid,c(hd,tok,opts opts0,body)) @
                      delay
                 else mtDM  --- stop resending
                 fi)  .
Passing time.
For managing the passing of time, we follow the Real-time Maude approach [olveczky08tacas]. The idea is that there is an earliest time at which some instantaneous rule will be enabled to fire, called the minimal time elapse (mte). If the mte is zero then the enabled rules should be applied until there are no more. If the mte is greater than zero then time passes by this amount, reaching a situation where some rules can fire at the current time (or the state is terminal). It is the job of the specification designer to ensure that only a finitely many rules can fire before time must pass. In the CoAP model, mte is greater than zero if all delays of messages in the network, or an w4Ack attribute, are greater than zero, or a message send is pending but the sndCtr is non-zero. The rule labelled tick formalizes the above. Here nz has sort NzNat (non-zero natural number). The function passTime decrements each message delay in the network and w4Ack attributes and each sndCtr by nz. It also decrements delays of messages in rspSntD attributes, modeling the decrease in lifetime of the identifier and token.
    crl[tick]:
    {conf} => {passTime(conf,nz,mt)}
    if nz := mte({conf}) .
3.3 Attack specification
Our specification of attacks against CoAP messaging consists of a specification of attack capabilities (sort Cap), a function doAttack, an attacker agent, and a rule for executing attacks, that invokes doAttack on a target message using an available capability. Attack models resulting from different choices of capability sets, and corresponding doAttack clauses, are discussed in Section 4. We note that the usual symbolic model matching patterns representing attacker ability to construct messages doesn’t work because there is no a priori expectation of what messages are expected.
An attacker agent has the same form as a device endpoint
      [aid  |  caps(acaps) attrs]
but with attacker specific attributes that include, caps(acaps), the capabilities attribute specifying what the attacker can do. The rule labelled attack specifies the semantics of an attack capability. It selects a target delayed message, dmsg, from the network input component, and a capability from the attackers capability set, and calls doAttack with the current attributes, the target message and the capability. The result is a replacement, dmsgs for dmsg and possibly updated attributes. For example, if the capability corresponds to drop then dmsgs is the empty set, and if the capability corresponds to replay after delay $n$ then dmsgs is dmsg dmsg1 where dmsg1 is dmsg with $n$ added to its delay.
    crl[attack]:
      [aid  |  caps(acaps acap) attrs]
       net(dmsgs0 dmsg,dmsgs1)
      [aid  |  attrs1]
      net(dmsgs0,dmsgs1 dmsgs)
     if toSend(dmsgs) attrs1 :=
        doAttack(attrs caps(acaps),dmsg,acap) .
3.4 CoAP Scenario Specification
A scenario is specified by an intial system configuration. This together with the CoAP rewrite theory determines a set of execution traces and the reachable states of the scenario. To support systematic testing of CoAP messaging and analysis of attacks we defined two functions to generate initial configurations for scenarios tCS2C (test Client Server 2 endpoints) and tCS3C (test Client Server 3 endpoints). These are used in the experiments validating the vulnerabilities discussed in [coap-attacks], described in Appendix 0.B. They were also used in experiments testing various CoAP messaging patterns and effects of dialecting.
In addition to having an empty network, initial configurations are characterized by the initial attributes of the endpoints. The function initDevAttrs defines the initial values of attributes that are the same for all endpoints in the scenarios considered. It has two arguments that are passed on to the function mkCoapConf the defines global configuration parameters
    op mkInitDevAttrs : Nat Nat -> Attrs .
    eq mkInitDevAttrs(mqd:Nat,w4ab:Nat) =
      w4Ack(mtDM)
      w4Rsp(mtM)
      rspSntD(mtDM)
      rspRcd(mtM)
      ctr(0)
      config(mkCoapConf(mqd:Nat,w4ab:Nat)) .
To experiment with different delays for sending application requests mkCoapConf has two arguments giving the values of the "msgQD" and "w4AckBd" configuration parameters. Recall that "msgQD" is the delay between message sends, and sending is blocked if there are more than "w4AckBd" confirmable messages awaiting acknowledgement.
    op mkCoapConf : Nat Nat -> CBnds .
    eq mkCoapConf(mqd:Nat,w4ab:Nat) =
      cb("ACK_TIMEOUT", 5)
      cb("ACK_RANDOM_FACTOR", 2)
      cb("MAX_RETRANSMIT", 1)
      cb("msgSD",2)
      cb("msgQD",mqd:Nat)
      cb("w4AckBd",w4ab:Nat)
      cb("ttl", 10) .
We define initial system configurations as follows.
Definition 1 (Initial CoAP configuration)
An initial CoAP system configuration sysI contains one network configuration element of the form
    net(mtDM,mtDM)
and a finite number of endpoint agents of the form
    [eid | sendReqs(amsgl:AMsgL) rsrcs(rbnds:RMap)  initDevAttrs(mqd:Nat,w4ab:Nat)]
An initial system configuration with attacker also has one attack agent of the form
     [ "eve" | kb(dmsgs) caps(caps:Caps)]
which we omit if caps:Caps is empty.
Initial System Configuration Generators.
The scenario generators use auxiliary functions to generate endpoint agents (mkDevC) and an optionally an attack agent (mkAtt).
The device agent generator mkDevC has arguments specifying its identifier (n), initial delay before sending messages (j), its list of application messages to send (amsgl), the initial state of any resources (rbnds), and two arguments passed on to mkInitDevAttrs.
    op mkDevC : Nat Nat AMsgL RMap Nat Nat -> Agent .
    ceq mkDevC(n,j,amsgl,rbnds,mqd:Nat,w4ab:Nat) =
      [ epid | sendReqs(amsgl) rsrcs(rbnds)
               sndCtr(j) mkInitDevAttrs(mqd:Nat,w4ab:Nat) ]
    if epid := "dev"  + string(n,10) .
The attack agent generator mkAtt has one argument, the capability set caps.
    op mkAtt : Caps -> Agent .
    eq mkAtt(caps) = ["eve" | kb(mtDM) caps(caps)] .
The function tCS2C has three groups of arguments, one group for each of two devices and one (singleton) group for the attacker. The resul is a client device, "dev0", with a list of application messages amsgl to send; a server device, "dev1", with resources rmap; and an attacker, "eve", with capabilities caps. The function simply applies the appropriate agent constructor to each group of arguments and adds an initial empty network state. If there are no attack capabilities (mtC), then no attack agent is generated.
****  2 endpoint configuration, with devconfig parameters
**** w4ab = n requires w4Ack to have <= n dmsgs
**** mqd = m requires delay of m before sending next request
    op tCS2C : Nat Nat AMsgL RMap Nat Nat
               Nat Nat AMsgL RMap Nat Nat
              Caps -> Conf .
    eq tCS2C(n0,j0,amsgl0,rbnds0,mqd0:Nat,w4ab0:Nat,
             n1,j1,amsgl1,rbnds1,mqd1:Nat,w4ab1:Nat,
            caps) =
       net(mtDM,mtDM)
       mkDevC(n0,j0,amsgl0,rbnds0,mqd0:Nat,w4ab0:Nat)
       mkDevC(n1,j1,amsgl1,rbnds1,mqd1:Nat,w4ab1:Nat)
       (if caps == mtC
        then mt
        else mkAtt(caps)
        fi)  .
The definition of tCS3C is similar, it just has one more group of device arguments and generates a third device.
A typical two endpoint scenario has one endpoint with some messages, amsgl0, to send (a client role) and one with initial resources rbnds1 (a server role) with message sending delay 5 and w4Ack bound of 0.
    op tCS : AMsgL RMap Caps -> Conf .
    eq tCS(amsgl0,rbnds1,caps) =
        tCS2C(0,0,amsgl0,mtR,5,0,
              1,1,nilAM,rbnds1,5,0,
             caps) .
A similar scenario but with two servers (with initial resources rbnds1 and rbnds2) is created by the function tCSS.
    op tCSS : AMsgL RMap RMap Caps -> Conf .
    eq tCSS(amsgl0,rbnds1,rbnds2,caps) =
        tCS3C(0,0,amsgl0,mtR,5,0,
              1,1,nilAM,rbnds1,5,0,
              2,1,nilAM,rbnds2,5,0,
             caps) .
Finally, we define application message constructors to use in creating specific scenarios. The functions mkGetC / mkGetN construct confirmable/non-confirmable GET messages. The arguments are a message identifier (id), the receiver (tgt), and the resource path (path).
    ****         Id     Tgt   Resource
    op mkGetC : String String String -> AMsg .
    op mkGetN : String String String -> AMsg .
    eq mkGetC(id,tgt,path) =
         amsg(id,tgt,"CON","GET",path,"",mtBody) .
    eq mkGetN(id,tgt,path) =
         amsg(id,tgt,"NON","GET",path,"",mtBody) .
Similarly the functions mkPutC / mkPutN construct confirmable/non-confirmable PUT messages with an additional argument, the value (val) to assign to the resource.
    ****         Id    Tgt    Resource  Val
    op mkPutC : String String String  String -> AMsg .
    op mkPutN : String String String  String -> AMsg .
    eq mkPutC(id,tgt,path,val) =
         amsg(id,tgt,"CON","PUT",path,"",b(val)) .
    eq mkPutN(id,tgt,path,val) =
         amsg(id,tgt,"NON","PUT",path,"",b(val)) .
The function mkDelN creates a non-confirmable DELETE message with the same arguments as the GET messages.
    ****         Id    Tgt    Resource
    op mkDelN : String String String -> AMsg .
    eq mkDelN(id,tgt,path) =
         amsg(id,tgt,"NON","DELETE",path,"",mtBody) .
In order to search for attacks based on order of events (for example, message receives) we introduce a Log configuration element.
    sort Log .   subsort Log < ConfElt .
A log is a list of log items specified as follows
    sorts LogItem LogItemL .
    subsort LogItem < LogItemL .
    op nilLI : -> LogItemL [ctor] .  --- the empty list
    op _;_ : LogItemL LogItemL -> LogItemL
            [ctor assoc id: nilLI] .  --- list concatenation
op log : LogItemL -> Log [ctor] .  --- the log entry
Receive of a request to PUT value val at resource path path by endpoint with identifier epid is represented by the log item
        rcvP(epid,path,val) .
Logging is currently only used by the rcvPut function that adds the attribute
  toLog(rcvP(epid,path,val))
to the returned attribute set. The receive rule crl[rcv] is augmented with the conditional clauses
    /\ conf1 := doLog(conf,devatts1)
    /\ devatts2 := clearToLog(devatts1)
to add any logitems from a toLog attribute to the log configuration element, and remove any toLog attribute from the device attributes.
To support defining properties based on ordering of events we define a function findRcvLI that given a log item list, a starting index, and a pattern (three string variables) returns a pair (sort LogItemIx) consisting of the first matching event on or after the starting index, together with the index of the event in the list.
    sort LogItemIx .
    op ‘{_‘,_‘} : LogItem Nat -> LogItemIx [ctor] .
    op findRcvLI : LogItemL Nat String String String ->  [LogItemIx] .
    op findRcvLIX : LogItemL String String String  Nat ->  [LogItemIx] .
    eq  findRcvLI(lil:LogItemL,n,epat,ppat,vpat)
        = findRcvLIX(nthCdr(lil:LogItemL,n),epat,ppat,vpat,n) .
    eq findRcvLIX(li:LogItem ; lil:LogItemL,epat,ppat,vpat,n)
      (if matchesLI(li:LogItem,epat,ppat,vpat)
       then {li:LogItem,n}
       else findRcvLIX(lil:LogItemL,epat,ppat,vpat,s n) fi) .
The nthCdr(lil:LogItemL,n) function returns the suffix of the logitem list lil:LogItemL beginning at the nth element (counting from 0). The logitem rcvP(epid,path,val) matches the pattern (epat,ppat,vpat) if corresponding elements of the triple match. A string matches a pattern string if the two are equal or the pattern is "". An error value results if no matching event is found.
For example to check that a request to "dev1" to unlock its door resource was followed by one locking that door we can use the property
    subLIL(c:Conf,rcvP("dev1","door","unlock") ;
                  rcvP("dev1","door","lock"))
where subLIL checks that there is a sublist of the configuration log element that matches the second argument logitem pattern list. It is defined using the findRcvLI function as follows.
    op subLIL : Conf LogItemL -> Bool .
    eq subLIL(c:Conf log(litl),plitl) = subLILX(litl,plitl,0) .
    op subLIXL : LogItemL LogItemL Nat -> Bool .
    ceq subLILX(litl,rcvP(epat,ppat,vpat) ; plitl,n) =
          subLILX(litl, plitl,n0) =
     if {rcvP(epid0,path0,val0), n0} :=
             findRcvLI(litl,n,epat,ppat,vpat)
    eq subLILX(litl,nilLI,n) = true .
    eq subLILX(litl,plitl,n) = false [owise] .
3.5 Sample scenarios.
To illustrate the use of the CoAP specification, and introduce some of the attacks of interest we define some small scenarios and show how to search for attacks. Attack models are discussed in more detail and formalized in Section 4.
Consider a scenario with client, "dev0", that requests server, "dev1", to unlock (PUTCDU), then lock the door (PUTCDL). In between the resource "sig" is set to "go" (PUTNSG). After a normal execution of this protocol, the door is locked at the end.
      dev0           dev1
      ----           lock
       o --PUTCDU-->  o
       o <--ack --    o
       o <--2.04--    o
       o --PUTNSG-->  o
       o <--2.01--    o
       o --PUTCDL-->  o
       o <--ack --    o
       O <--2.04--    o
We use the function tCS and the application message constructors described above to create the initial configuration for this scenario:
    iSys0 = {tCS(mkPutC("putCDU","dev1","door","unlock") ;
                 mkPutN("putNSG","dev1","sig","go") ;
                 mkPutC("putCDL","dev1","door","lock"),
                 rb("door","lock"),mtC)} .
Now we ask: Can the scenario (with no attacker) end with the door unlocked if the attacker has no capabilities? This is done by searching for a terminal state reachable from iSys0 in which the "dev1" "door" resource has value "unlock".
    search [1] iSys0 =>! {c:Conf} such that
       checkRsrc(c:Conf,"dev1","door","unlock") .
There is no solution. Next we ask if the attacker who can drop a message can force the scenario end with the door unlocked? iSys1 adds the drop capability to the configuration iSys0.
    iSys1 = {tCS(mkPutC("putNDU","dev1","door","unlock") ;
                 mkPutN("putNSG","dev1","sig","go") ;
                 mkPutC("putNDL","dev1","door","lock"),
            rb("door","lock"),drop)}
We search for a final state where the "door" resource of "dev1" has value "unlock".
    search [1] iSys1 =>! {c:Conf} such that
        checkRsrc(c:Conf,"dev1","door","unlock") .
There is no solution because even if the attacker drops the lock request, since the request is confirmable, the client will retry. If the attacker can drop two messages then it can force termination with the door unlocked [not shown].
Suppose the attacker can replay a message (i.e. make a copy and delay it) as in scenario iSys2.
    iSys2 = {tCS(mkPutC("putNDU","dev1","door","unlock") ;
                 mkPutN("putNSG","dev1","sig","go") ;
                 mkPutC("putNDL","dev1","door","lock"),
              rb("door","lock"),replay(10))}
Can the attacker cause an unlock to follow a lock request? We search for an execution in which a door lock request precedes an unlock request. (We could also require that the unlock request was the last door request.)
    search iSys2 =>! {c:Conf} such that
        subLIL(c:Conf, rcvP("dev1","door","lock") ;
                       rcvP("dev1","door","unlock") ) .
There are two solutions. The log from the second solution follows.
  log(rcvP("dev1","door","unlock") ; rcvP("dev1","sig","go") ;
      rcvP("dev1","door","lock") ; rcvP("dev1","door","unlock"))
Now, consider a situation where there are two servers ("dev1" and "dev2") with door resources. The client wants the state of the door at "dev1". Can the attacker reroute the "GET" request so that the client receives the state of the door at "dev2" instead?
Here we use and instance of a general attack capability
   mc(tgt,src,b:Bool,act(tpat,spat,d:Nat))
This capability applies to messages to tgt from src. It changes the target and source according to tpat and spat ("" means no change) and adds d:Nat to the delay. If the flag b:Bool is true the edited message replaces the original, while if the flag b:Bool is false the original message is unchanged and the edit is a copy.
      dev0         eve           dev1        dev2
      ----         ---          unlock       lock
       o -GETN(1)-> @ [-GETN(1)->] ?
       |            o               -GETN(2)->o
       | <-lock(1)- @ <---lock(2)--            o
       |          [<--unlock(1)--] o
       o <-oneof{lock,unlock}-
iSys3a is the scenario for the case when the attacker modifies a message in transit. Note that the attacker must redirect both the request (changing the target) and the response (changing the source).
    iSys3a = tCSS(mkGetN("getN","dev1" ,"door"),
                  rb("door","unlock"),rb("door","lock"),
                 mc("dev1","dev0",true,act("dev2","dev0",0))
                 mc("dev0","dev2",true,act("dev0","dev1",0))) .
To check for an attack we search for a final state in which the door resource of "dev1" is unlocked, but "dev0" receives a response (apparently) from "dev1" with the value "lock".
    search iSys3a =>! {c:Conf} such that
        checkRsrc(c:Conf,"dev1","door","unlock") and
        hasGetRsp(c:Conf,"dev0","dev1","getN","lock") .
There are 4 solutions of 33 states visited. The log is empty because the model currently does not currently log "GET" requests.
Suppose the attacker can not change messages in transit, but can make and edit copies. This is modelled by setting the flag in the capabilities to false, as in the scenario iSys3r.
    iSys3r = tCSS(mkGetN("getN","dev1" ,"door"),
                  rb("door","unlock"),rb("door","lock"),
                  mc("dev1","dev0",false,act("dev2","dev0",0))
                  mc("dev0","dev2",false,act("dev0","dev1",0))) .
We repeat the search for attacks starting with iSys3r using this more restricted attack.
    search iSys3r =>! {c:Conf} such that
        checkRsrc(c:Conf,"dev1","door","unlock") and
        hasGetRsp(c:Conf,"dev0","dev1","getN","lock") .
There are also 4 solutions. In this case 109 states were visited, indicating that the probability of attacker success is less when it can only make copies and not change the original message in transit.
4 Attack Models
We consider two main classes of attacker: active and reactive. The active attacker models the attacker envisioned in [coap-attacks]. Using the formal representation of an active attacker we replicated versions of the attacks postulated in [coap-attacks] (see Appendix 0.B). These attack capabilities (except for the rerouting extension, A1, see below) simply amplify the effects of an unreliable transport, dialecting can not mitigate such attacks, and applications need to deal with the unreliability of the network independently of attackers.
The reactive attacker can not modify messages in transit (can not drop or delay these messages). But it can observe, make copies of messages in transit, and transmit modified copies (redirecting, resending with delay). These capabilities violate the expected network guarantee that if an endpoint, ep1 receives a message with source ep0, then there is a unique previous event in which endpoint ep0 sent that message. They are also capabilities that can be mitigated by dialecting as we show in Section 8.
The drop capability of scenario iSys1 of Section 3.5 is an example of an active attack. The replay capability of scenario iSys2 can be used by active or reactive attackers. The flag in the redirection capability of scenarios iSys3a and iSys3r determines whether the attack is active (flag is true) or reactive (flag is false). This is formalized in Section 4.4
4.1 Active attacker
A simple active attacker (A0) has a finite set of single actions from the capability set $\mathtt{CapsA0}=\{\mathtt{drop},\mathtt{delay(n)}\}$, where drop (also called block) prevents a message from being delivered, and delay(n) delays delivery of a message by $n$ time units. $\mathtt{CapsA1}$ extends $\mathtt{CapsA1}$ with two additional capabilities $\{\mathtt{replay(n)},\mathtt{reroute(tpat,spat,n)}\}$. Applying replay(n) makes a copy of the message and transmits it with $n$ added to its delay, while reroute(tpat,spat,n) modifies the message target and source according to tpat and spat respectively and adds $n$ to the delay. The pattern tpat is either a wild card (the empty identifier) which leaves the target unchanged, or a specific endpoint identifier that replaces the original. Similarly for the action of spat on the message source.
Informally, these attack capacities can be used to achieve the following basic effects on CoAP messaging.
  • drop can be used to prevent a request (for some action) from being received, or to prevent a response to some request. In the response case this leaves the client is a state of not knowing the requested data value, or if the requested action happened. Note that due to CoAPs reliability mechanism, if the target message is confirmable, the drop attack will need to be applied to all retries to succeed.
  • The delay of a request, say DM1 from a sequence DM1 ... DM2 until after DM2 can undo the effect of DM2 (door unlock after door lock), or could cause DM2 to happen in the wrong context, for example without lights turned on.
  • The delay of a response, say to DM1 delivered as response to a later (blocked) request DM2 could cause the requesting client to receive out-of-date data, or to think the action requested by DM2 happened, when only the action of DM1 happened. If DM1, DM2 are GET requests for different resources, say r1 and r2, the client may interpret the response received as the value of r2 when it is the value of r1. We note that this only happens with weak implementation of the token mechanism used to link requests and responses.
  • Rerouting a request from ep0 to ep1 to a request from ep0 to ep2 and (un)rerouting the response from ep2 to be from ep1 causes the client ep0 to interpret the response as the value of requested resource at ep1 when it is the value at ep2.
The capability set CapsA0 is sufficient to demonstrate all the vulnerabilities studied in Appendix 0.B, except for the modified final example that uses rerouting.
4.2 Reactive attacker
A reactive attacker can not modify or block a message in transit. It can see/copy a message, change the source/target or the delay amount and transmit the result.
A simple reactive attacker has capabilities in the set $\mathtt{CapsR0}$ that allows the attacker to make (and edit) one copy of a target message. A given attacker instance is limited to a finite number of attack attempts.
We also consider a multi action reactive attacker, with capabilities in the set $\mathtt{CapsR1}$. Using a multi-action, an attacker can make multiple copies per message, optionally adding to the delay and/or modifying the message source or target. We allow the attacker to restrict attention to a given target-source pattern. This mainly improves the attacker efficiency (reduces the search space) but adds no power at the granularity we consider. As for the simple reactive attacker, each attacker instance is limited to a finite set of multi-action capabilities. In a little more detail a multi-action capability consists of a condition match(tpat,spat) constraining the set of messages to be attacked, and a finite set of attack actions act(tpat,spat,n). As before, (tpat,spat) can be wild cards (represented by the empty string) or specific endpoint identifiers. In the match case a wild card matches any identifier, while specific identifiers require equality to match. In the case of act, a wild card means the corresponding target or source is not changed, while a specific identifier causes the corresponding target or source to be replaced by the pattern element. The copy is further delayed by the amount given by the number $n$.
The simple reactive attacker is a special case of the multi-action reactive attacker, where the condition is restricted to match("","") and the capset applied to a message is a singleton. Although a reactive attacker can’t block or delay messages in transit, many of the active attack effects can be achieved.
Here are examples of what a reactive attacker can do.
  • R1. Undo/revert an action. Suppose messages M1, M2 PUT different values for the same resource. The attacker copies M1 and sends the copy with sufficient delay to arrive after M2, thus overriding the effect of M2. This emulates some of the reordering capability of the active attacker, leaving the resource in a state other that what the client planned. It could for example leave a process running that should have stopped. If the client then starts a process by PUTting a different resource in the state "on" after sending M2, the new process may run in an unexpected context.
  • R2. Violate ordering or concurrency constraints. The attacker observes a message M1 from ep0 to ep1. It makes copies redirected to ep2 (ep3 …). A compliant CoAP endpoint ep0 will ignore responses from ep2 (ep3 …). But in the case of requests causing actions, there will be unexpected actions. For example, if ep0 is activating ep1, ep2, …, in sequence then ep2, …, will be acting out of sequence, concurrently with ep1, which could lead to undesired consequences.
  • R3. Duplication of a process. Suppose as above ep0 is coordinating a process by activating ep1, …, epk in sequence, using messages M1, …, Mk. Suppose also that ep0x, ep1x, …, epkx are a functionally equivalent set of endpoints, say in a different location. Then the attacker can copy each message and redirect to ep1x, …epkx.
  • R4. Spoofing (Redirect GET request/response). The attacker observes a message M1 from ep0 to ep1. It makes a copy redirected to ep2. A compliant CoAP endpoint will reject the response from ep2 (wrong server). If the attacker also makes a copy of the response, rerouting it to appear to be from ep1 then ep0 will accept which ever response arrives first. If M1 is a GET request, then the client may receive the wrong value for the requested resource.
R1 only needs one capability instance with one action. R2 only needs one capability instance but with 2 or more actions. R3, R4 need multiple capability instances, each one needs only a single action.
4.3 More possibilities
There are additional, stronger, attack models that could be considered in the future. One is a reactive attacker with memory. Here the attacker sees all messages and can remember some. Then at some point the attacker can make edited copies of some of the messages in its knowledge base and transmit them, for example to rerun an application sequence either with the original set of endpoints at a later time, or a separate set (with the same behavior) at some point. The attacker may use observation of a message occurrence to trigger actions, or time elapse.
A creative reactive attacker can additionally create new messages and receive responses, i.e. becoming a rogue endpoint.
4.4 Attack-model specification.
As part of the CoAP specification, we specified a general attack framework where an attacker is endowed with a finite set of capabilities. In this section we specify specific capabilities corresponding to the active and reactive attack models discussed above. We define capability constructors and define their semantics by giving equations for the doAttack function for each capability. Different models are obtained by constraining the capability attribute of the attacker agent.
We define a generic capability construction, mc(tpat,spat,active?,acaps). tpat,spat are target, source patterns used to restrict the set of messages to be ‘attacked’. acapsis a possibly empty set of actions. Each action, act(tpat1,spat1,d), causes a copy of the matched message to be made, transformed, and transmitted. tpat1,spat1 are target, source patterns used to determine he target and source of the transformed message. If a pattern is "" the original target or source is used, otherwise the pattern string is used. The delay d is added to the delay of the original message. If the boolean active? is true, this is an active capability, and the original message is removed from the network, only the copies generated by the actions are transmitted. If active? is false the capability is reactive and the original message is left in the network to be transmitted as usual. This is formalized by the following definitions.
    ****      tpat   spat  delay
    op act : String String Nat -> Cap .
    ****     tpat  spat  active? actions
    op mc : String String Bool Caps -> Cap .
The function doAttack defines the semantics of each attack capability. It is given the attack agent attributes, including its knowledge base of messages seen, the attack target message, dmsg, and a multi capability. If the target message matches the selection pattern (using pmatch) then the toSend attribute is computed and returned along with the knowledge base with the attacked message added. The toSend attribute includes the attacked message iff the flag b is false. It also includes the results of applyCaps(dmsg,acaps,mtDM), that gives the semantics of each action.
    ****                tgt  act   result
    op doAttack : Attrs DMsg Cap -> Attrs .
    ceq doAttack(attrs kb(dmsgs),dmsg,mc(tpat,spat,b,acaps))
           = toSend(dmsgs1) kb(dmsgs dmsg) attrs
     if msg @ n := dmsg
     /\ pmatch(getTgt(msg),tpat)
     /\ pmatch(getSrc(msg),spat)
     /\ dmsgs0 := applyCaps(dmsg,acaps,mtDM)
     /\ dmsgs1 := dmsgs0 (if b then mtDM else dmsg fi) .
If the match fails doAttack returns a result that does not match the expectation of the attack rule, and the rule does not fire.
The function applyCaps makes a transformed copy of the attacked message for each action in acaps. The transformation sets the message target and source according to the action patterns, and increments the delay,
**** multi cap attack
    op applyCaps : DMsg Caps DMsgS -> DMsgS .
    eq applyCaps(dmsg,mtC,dmsgs) = dmsgs .
    eq applyCaps(msg @ n,act(tpat,spat,n0) acaps,dmsgs) =
          applyCaps(msg @ n,acaps, dmsgs
          (setTgtSrc(msg,tpat,spat) @ (n + n0)) ) .
For early experiments and readability we defined a number of specific (active) attack capacities. We list them here with their definitions as instances of the mc construction.
  • drop = mc("","",true,mtC) — drop the target message
  • delay(n) = mc("","",true,act("","",n)) — add n to the target message delay
  • divert(dst0,dst1) = mc(dst0,"",true,act(dst1,"")) — reroute a target message with destination dst0 to dst1
  • undivert(src0,src1) = mc("",src0,true,act("",src1,0)) — make a target message from src0 appear to be from src1
  • replay(n)= mc("","",false,act("","",n)) — replay the target message after delay n (the original message is delivered as expected)
5 Example reactive attacks
This section describes experiments using scenarios illustrating the 4 classes of reactive attack discussed in Section 4.
5.1 R1. Redo/revert an action
The R1 scenario is a variant of the attacks on CoAP vulnerabilities Figures 3,4,5,6 (see Section 0.B). The basic requirement is that the door remains unlocked. The attacker aims for the door to be locked at the end, or before the optional signal message.
        dev0        eve        dev1
        ----        ---        unlock
         o -PUTNDL-> ----------->
                    @ act("","",5/10)
         o <-------  <-2.04--    o
         o -PUTNDU-> ----------> o
         o <-------  <-2.04--    o
                    --- PUTNDL-> o  --- attack
    optional
         o -PUTNSO->  ------>    o
         o <-------  <-2.01--    o
The function raR1 generates instances of this scenario. It is parameterized by the delay between application message sends (mqd:Nat), the bound on pending acknowledgements that blocks sending an application message (w4b:Nat), the delay to be added to an replayed message (d:Nat), and a boolean (nso:Bool) specifying whether the optional "signal" message should be sent.
    op raR1 : Nat Nat Nat Bool -> Conf .
    eq raR1(mqd:Nat,w4b:Nat,d:Nat,nso:Bool) =
     tCS2C(0,0,mkPutN("putNDL","dev1","door","lock") ;
               mkPutN("putNDU","dev1","door","unlock") ;
             (if nso:Bool
              then mkPutN("putNS","dev1","signal","on")
              else nilAM fi) ,
              mtR,mqd:Nat,w4b:Nat,
          1,1,nilAM,rb("door","unlock")rb("sig","off"),2,0,
          mc("dev1","dev0",false, act("","",d:Nat))) .
Can the attacker cause the execution to end with the door locked? To check, search for a terminal state in which "dev1" has resource "door" with value "lock".
    search {raR1(5,0,10,false)} =>! {c:Conf} such that
        checkRsrc(c:Conf,"dev1","door","lock") .
There are 2 solutions in 101 states visited.
To check that the attacker can shut the door before a later event, we turn on the signal option and search for a final state where the lock event happened before the signal event.
    search {raR1(5,0,10,true)} =>! {c:Conf} such that
        checkRsrc(c:Conf,"dev1","door","lock") and
        subLIL(c:Conf,rcvP("dev1","door","unlock") ;
                      rcvP("dev1","door","lock") ;
                      rcvP("dev1","signal","")) .
There are 2 solutions from 231 states visited. The log of the second solution is the following.
    log(rcvP("dev1","door","lock") ;
        rcvP("dev1", "door", "unlock") ;
        rcvP("dev1", "door", "lock") ;
        rcvP("dev1", "signal", "on"))
If the attacker uses a delay smaller than 10 or larger than 10 there is no attack (search yields no solution).
5.2 R2. Violating ordering or concurrency constraints
In this example, a sequence of $n$ tasks is to be executed one after the other. The protocol consists of a controller (client endpoint) and n servers that execute the tasks. A task is initiated when the server receives a PUT "sig" "on" request and terminates when he server receives a PUT "sig" "off" request.
    dev0        eve        dev1  ... devk
    ----        ---        off
     o -PUTNon  ----------> o
                 @(cc dev2)> ----->
     o <-------  <-2.04--   o
     o -PUTNoff  ----------> o
     o <-------  <-2.04--    o
     o -PUTNon-> ---------->         o
     o <-------  <-2.04---------     o
     o -PUTNoff-> -----------------> o
     o <-------  <-2.04---------     o
The function iSysX generates instances of the above scenario. n is the number of servers, d the task duration, and caps the attacker capabilities.
    op iSysX : Nat Nat Caps -> Sys .
    eq iSysX(n,d,caps) = {CnS(n,mkSigAMs(n,d,nilAM),rb("sig","off"),5,0)
      mkAtt(caps) } .
It uses the function CnS(n,amsgl,rbnds,mqd,w4ab) that generates a configuration with one client and n servers. The client’s application message list is amsgl and each server has initial resources given by the RMap rbnds. The client delay between message sends is mqd and it can only send a new message if there are no more than w4ab confirmable messages awaiting an ACK.
The function mkSigAMs(n,d,nilAM) generates the list of application messages for the client to send to control the execution. The list consists of pairs
   mkPutN("putN","dev" + string(j,10),"sig","on") ;
   mkPutN("putN","dev" + string(j,10),"sig","off") .
for $j\in[1,n]$ separated by a delay of d.
We consider three levels of attacker capability. Level $j$ (caps-j) looks for messages from "dev0" to "dev1" and immediately (additional delay 0) sends a copy to "dev2""dev1+j".
    ops caps-1 caps-2 caps-3 : -> Caps .
    eq caps-1 = mc("dev1","dev0",false,act("dev2","",0)) .
    eq caps-2 = mc("dev1","dev0",false,
                   act("dev2","",0) act("dev3","",0)) .
    eq caps-3 = mc("dev1","dev0",false,
                   act("dev2","",0) act("dev3","",0)
                   act("dev4","",0)) .
We consider two attacks that violate the property that the tasks do not interleave, by starting a task on "dev-k+j" before the task on "dev-k" completes. One check for violation is that the number of devices with resource binding rb("sig","on") is great than 1. The function epswrb(conf,rbnds) returns the number of servers with resource map containing a match for rbnds is used for this check. Alternately we can check for executions in which "devj" receives an "on" signal, and "devj+1" receives the "on" signal before "devj" receives the "off" signal. This is done using the subLI function. Using caps-1 can the attacker can cause at least 2 simultaneous tasks executions?
    search iSysX(3,0,caps-1) =>+ sys:Sys such that
        size(epswrb(sys:Sys,rb("sig","on"))) > 1 .
There are 132 solutions from 767 states visited.
Raising the bar we ask if using caps-2 can the attacker cause 3 or more tasks to execution simultaneously.
    search  iSysX(3,0,caps-2) =>+ sys:Sys such that
        size(epswrb(sys:Sys,rb("sig","on"))) > 2 .
There are 182 solutions among 721 states visited.
We can also ask if the attacker can cause dev2 to start before dev1 finishes.
    **** attack 3 dev2 starts before dev1 finishes
    search iSysX(3,0,caps-2) =>+ {c:Conf} such that
        subLIL(c:Conf, rcvP("dev1","sig","on") ;
                       rcvP("dev2","sig","on") ;
                       rcvP("dev1","sig","off")) .
There are 342 solutions from 3598 states visited. The log for the last solution is
   log(rcvP("dev3", "sig", "on") ; rcvP("dev1", "sig", "on") ;
    rcvP( "dev2", "sig", "on") ; rcvP("dev1", "sig", "off") ;
    rcvP("dev2", "sig", "on") ; rcvP("dev2", "sig", "off") ;
    rcvP("dev3", "sig", "on") ; rcvP( "dev3", "sig", "off")).
Finally, we ask if the attacker can cause dev2 to start before dev1 finishes and dev3 to start before dev2 finishes.
    search iSysX(3,0,caps-2) =>+ {c:Conf} such that
        subLIL(c:Conf, rcvP("dev1","sig","on") ;
                       rcvP("dev2","sig","on") ;
                       rcvP("dev1","sig","off")) and
        subLIL(c:Conf,rcvP("dev2","sig","on") ;
                      rcvP("dev3","sig","on") ;
                      rcvP("dev2","sig","off")) .
There are 62 solutions among 3594 states visited. The log for the last solution follows.
  log(rcvP("dev1", "sig", "on") ; rcvP("dev2", "sig", "on") ;
  rcvP( "dev3", "sig", "on") ; rcvP("dev1", "sig", "off") ;
  rcvP("dev2", "sig", "on") ; rcvP("dev2", "sig", "off") ;
  rcvP("dev3", "sig", "on") ; rcvP( "dev3", "sig", "off"))
5.3 R3 Duplication of a process.
IoT systems may be replicated, for example running the same protocol in multiple smart rooms, or in a manufacturing setting multiple instances of a manufacturing step.
    dev0        eve        dev1 dev2 dev3 dev4 dev5 dev6
     o -PUTNgo   @ ---------> o
                 o -(cc dev4)-------------> o
     o <-------  <-2.04--   o
                                 <-2.04--   o
     o -PUTNgo-> @-------------->  o
                 o -(cc dev5) ---------------->  o
     o <-------  <-2.04---------  o
                                       <-2.04--  o
     o -PUTNgo   @ -------------------> o
                 o -(cc dev6)----------------------> o
     o    <-------  <-2.04---------    o
                                            <-2.04-- o
Here the label (cc dev4) indicates the attacker coping a message to a different device.
The function iSysY generates example scenarios with one client (controller) and 2 sets of n servers. The client will send a PUT message to set the "sig" resource "go" to each server in the first set of servers. Each server initially has resource map rb("sig","off").
    op iSysY : Nat  Caps -> Sys .
    eq iSysY(n,caps) =
          {CnS(2 * n,mkGoAMs(n,nilAM),rb("sig","off"),5,0)
           mkAtt(caps)  log(nilLI)} .
We consider scenarios with $n=2$ and $n=3$ and ask whether the attacker can drive a the second set of devices, either concurrently or following the first. The attacker capabilities for the two scenario sizes are defined as follows, parameterized by the amount to delay message copies.
    ops caps2-2 caps3-3 : Nat -> Caps .
    eq caps2-2(d) =
      mc("dev1","dev0",act("dev3","dev0",d))
      mc("dev2","dev0",act("dev4","dev0",d)) .
    eq caps3-3(d) =
      mc("dev1","dev0",act("dev4","dev0",d))
      mc("dev2","dev0",act("dev5","dev0",d))
      mc("dev3","dev0",act("dev6","dev0",d)) .
The four search commands, results (with witness log) are as follows.
     *** search for duplicate 2 server process running concurrently
    search iSysY(2,caps2-2(0)) =>! {c:Conf} such that
      subLIL(c:Conf, rcvP("dev3","sig","") ; rcvP("dev4","sig","")) and
       subLIL(c:Conf, rcvP("dev3","sig","") ; rcvP("dev2","sig","")) .
There are 16 solutions among 534 states visited.
    log(rcvP("dev3", "sig", "go") ; rcvP("dev1", "sig", "go") ;
        rcvP( "dev4", "sig", "go") ; rcvP("dev2", "sig", "go"))
    *** search for duplicate 2 server process the second after the first
    search iSysY(2,caps2-2(15)) =>! {c:Conf} such that
      subLIL(c:Conf,rcvP("dev3","sig","") ;  rcvP("dev4","sig",""))
      and subLIL(c:Conf,rcvP("dev2","sig","") ; rcvP("dev3","sig","")) .
There are 4 solutions among 179 states visited.
    log(rcvP("dev1", "sig", "go") ; rcvP("dev2", "sig", "go") ;
        rcvP( "dev3", "sig", "go") ; rcvP("dev4", "sig", "go"))
     **** search for duplicate 3 server process the second after the first
    search iSysY(3,caps3-3(0)) =>! {c:Conf} such that
      subLIL(c:Conf, rcvP("dev4","sig","") ;
        rcvP("dev5","sig","") ; rcvP("dev6","sig","")) and
      subLIL(c:Conf, rcvP("dev1","sig","") ;
        rcvP("dev4","sig","") ; rcvP("dev2","sig","")) and
      subLIL(c:Conf, rcvP("dev2","sig","") ;
        rcvP("dev5","sig","") ; rcvP("dev3","sig","") ;
        rcvP("dev6","sig","")) .
There are 8 solutions among 2845 states visited.
    log(rcvP("dev1", "sig", "go") ; rcvP("dev4", "sig", "go") ;
        rcvP( "dev2", "sig", "go") ; rcvP("dev5", "sig", "go") ;
        rcvP("dev3", "sig", "go") ; rcvP("dev6", "sig", "go"))
    **** search for duplicate 3 server process the second after the first
    search iSysY(3,caps3-3(15)) =>! {c:Conf} such that
        subLIL(c:Conf, rcvP("dev4","sig","") ;
            rcvP("dev5","sig","") ; rcvP("dev6","sig","")) and
        subLIL(c:Conf, rcvP("dev3","sig","") ; rcvP("dev4","sig","")) .
There are 8 solutions among 683 states visited. The log of the last solution is the following.
   log(rcvP("dev1", "sig", "go") ; rcvP("dev2", "sig", "go") ;
       rcvP( "dev3", "sig", "go") ; rcvP("dev4", "sig", "go") ;
       rcvP("dev5", "sig", "go") ; rcvP("dev6", "sig", "go"))
5.4 R4. Spoofing (Redirect GET request/response)
Although a reactive attacker can not redirect a message in transit, as the active attacker can (see Section 0.B.6), it can make a copy and direct it to a different server. The reactive attacker also has to make a copy of the response from that server to appear to come from the original server. Thus the client will receive 3 responses. They can arrive in any order. The client will ignore the one from the alternate server, and its a race between the two messages apparently from the intended server.
        dev0      eve       dev1     dev2
        ----      ---      unlock    lock
      o -GETN(1)-->@------->  o
      |            o copy(cc dev2)  -> o
      o   <---unlock---      o
      o            @    <---lock---    o
      o <- (cc from dev1)
The function iSySZ(mqd:Nat,w4b:Nat) generates instances of the redirection attack scenario. It is parameterized by the usual congestion control arguments.
    op iSySZ : Nat  Nat -> Sys .
    eq iSySZ(mqd:Nat,w4b:Nat) =
         {tCS3C(0,0,mkGetN("getN0","dev1","door"),
                   mtR,mqd:Nat,w4b:Nat,
              1,1,nilAM,rb("door","unlock"),5,0,
              2,2,nilAM,rb("door","lock"),5,0,
              mc("dev1","dev0",false,act("dev2","",0))
              mc("dev0","dev2",false,act("","dev1",0)) )
              log(nilLI) }  .
To demonstrate attacks we search for final states in which the client has received a "lock" response to the GET request while "dev1" has the "door" resource bound to "unlock".
    search iSySZ(5,0) =>! {c:Conf} such that
      hasGetRsp(c:Conf,"dev0","dev1","getN0","lock") and
      checkRsrc(c:Conf,"dev1","door","unlock") .
There are 4 solutions among 109 states visited.
6 Dialect functions
Dialects using a moving target strategy must synchronize on the choice of lingo. In the case of reliable transport this is often done using time synchronization.  [^2] Even with reliable transport, attackers may drop or reorder messages, which can complicate lingo synchronization.
For protocols, like CoAP, running on unreliable transport, the notion of current lingo (or lingo parameters) or lingo (parameters) being no longer valid, seems problematic. In principle there is no bound on the delay of a message, and missing messages or out of order messages which are common, make guaranteeing that a receiver can determine the correct lingo to decode a dialected message a challenge. Failure would mean dropping a message that should have been delivered.
We take inspiration from security protocol designs that use message counters (one counter for each communication pair and direction) for replay prevention and/or detection. These counters appear in the clear in messages making synchronization simpler. We have to be careful that these counters don’t leak enough information to be useful to an attacker.
For the CoAP dialect specification we propose a scheme of three functions:
[…]
  • $f_{1}:\mathtt{String}\times\mathtt{Content}\times\mathtt{Nat}\longrightarrow\texttt{DCBits}$ – the obfuscator/encoder. The first argument is the source of randomness, the second argument is a message content and the third an index (a natural number).
  • $f_{2}:\mathtt{String}\times(\texttt{DCBits}\times\mathtt{Nat})\leadsto\texttt{ContentNat}$ – the de-obfuscator/decoder. The first argument is again the source of randomness and the second argument is a pair consisting of the encoded message content and the lingo index.
We can think of a dialect given by the above three functions as having a single parameterized lingo, or as having a family of lingos indexed by the third argument to $f1$.
There are two important properties of these functions. First, $f_{2}$ recovers the original content encoded by $f_{1}$, using random generator $g$
$f_{2}(grand,\{f_{1}(grand,\{\mathtt{content},ix\}\}),ix)=\{\mathtt{content},ix\}.$
[…]
$f_{2}(grand,\{\mathtt{dcbits1},ix\})=\{\mathtt{content1},ix\}.$
and
$f_{2}(grand,\{\mathtt{dcbits},ix1\})==\{\mathtt{content1},ix1\}.$
with $\mathtt{content}\neq\mathtt{content1}$ and $\mathtt{content1}$ is recognized as ill-formed. Alternately, $f_{2}$ could directly return an indication of failure, but only if the input is a correctly transformed content.
[…]
An alternative for generating the sequence of random strings is $g^{ix(}seed,k)$ $seed_{0}=g(seed,k,0)$ and $seed_{j+1}=g(seed_{j},k,0)$. If the attacker learns the value of $g(seed_{j},k,0)$ (for example if $f_{1}$ uses xor and the attacker can guess the content of some message) then the attacker can compute $g(seed_{j+n},k,0)$.
[…]
7 Specification of a CoAP dialect wrapper
To specify a dialect wrapper for CoAP messaging we need to specify the data type of dialected messages, the three functions described in Section 6, the structure of wrapped CoAP agents (subsection 7.1), and the rules for sending and receiving messages by the wrapped agents (Subsection 7.2). Towards dialects as a theory transformation, in subsection 7.3 we define operations D and UD. D takes an inital system configuration and produces the corresponding dialected configuration. UD extracts the underlying CoAP system configuration from a dialected configuration. Using these functions, rewrite and search commands for a CoAP scenario can be automatically transformed to rewrite and search commands for the corresponding dialected scenario (Subsection 7.4).
7.1 Dialect messages and configurations
A dialected message has a target and a source (sort String as for normal messages) and a dialect encoded content (sort DContent). The message constructor is overloaded to construct dialected messages.
    op m : String String DContent -> Msg .
A term of sort DContent is a pair constructed by the operator dc
    op dc : DCBits Nat -> DContent [ctor] .
where the sort DCBits is an opaque sort whose structure is not further specified. We also define the sort ContentNat to be pairs constructed from ordinary content and a natural number.
    op ‘{_‘,_‘} : Content Nat -> ContentNat [ctor] .
The three dialect functions of Section 6 are specified as follows:
    op g : String Nat Nat -> String .
    op f1 : String ContentNat -> DCBits .
    op f2 : String DContent  ~> ContentNat .
    eq f2(g(rand,k,ix),dc(f1(g(rand,k,ix),{content,ix}),ix))
          = {content,ix} .
[…]
As described in [GEMM-2023esorics] we represent a dialect wrapper by a meta-agent that has the same identifier as the wrapped agent, a conf attribute that encapsulates the base agent and its network stub, and additional attributes for managing dialect transformations.
    *** dialected endpoint
    [eid | conf([eid | devattrs] net), ddevattrs]
We refer to the encapsulated network as the internal network and the network at the level of meta entities as the external network.
The additional attributes for a dialect meta endpoint include:
  • used(umap) — a map from endpoint ids to sets of lingo indices already received from the identified endpoint.
  • toRcv(dmsgs) — holds the result of decoding– the original message or the empty (delayed) message set, if decoding fails
  • seedTo(eid,str) – the shared secret seed for sending to eid
  • seedFr(eid,str) – the shared secret seed for receiving from eid
  • ixCtr(eid,nat) — the counter for generating lingo indices for messages to endpoint named eid
  • randSize(k) – the size of generated (pseudo) random strings
7.2 Dialect Rules
There are two rules to specify dialect behavior. The rule labeled ddevsend handles applying a dialect lingo to a message sent by an endpoint and putting it in the global network. It selects a message in the internal network sent by the wrapped entity, calls applyDialect, and puts the resulting dialected message in the external network.
    crl[ddevsend]:
      [epid | conf([epid  | devatts ] net(dmsgs0,dmsgs1)) dattrs ]
      net(ddmsgs0,ddmsgs1)
      [epid | conf([epid  | devatts ] net(mtDM,dmsgs) ) dattrs1 ]
      net(ddmsgs0 ddmsg,ddmsgs1)
    if dmsgs (msg @ n) := dmsgs0 dmsgs1
    /\ getSrc(msg) = epid
    /\ toSend(ddmsg) dattrs1 := applyDialect(dattrs,msg @ n) .
The function applyDialect looks up seed, random size, and current index values in the attribute set dattrs. It computes the generated string grand using the dialect function g, and then applies the encoding function f1 to this string, the message content and the index and constructs the dialected message msgd to be sent on the external network. It also increments the index counter for the message target
    ceq applyDialect(dattrs,msg @ n) =
         incIxCtr(dattrs,dst,1) toSend(msgd @ n)
    if m(dst,src,content) := msg
    /\ rand := getSeedTo(dattrs,dst)
    /\ rsize := getRandSize(dattrs)
    /\ ix := getIxCtr(dattrs,dst)
    /\ grand := g(rand,rsize,ix)
    /\ dcbits := f1(grand,{content,ix})
    /\ msgd := m(dst,src,dc(dcbits,ix)) .
The rule labeled ddevrcv handles receipt of a dialected message. It selects a message with 0 delay and target the wrapped entity an calls decodeDialect. If decoding produced a message, the rule calls the rcvMsg directly since the result may include a logging attribute that needs the full configuration to process, which the base CoAP receive rule would not have.
    crl[ddevrcv]:
    {[epid | conf([epid | devatts ] net(dmsgs0,dmsgs1))
             dattrs ]
      net(ddmsgs0,ddmsgs1 msgd @ 0) conf}
    {[epid | conf([epid | devatts2 ]
                  net(dmsgs0 dmsgs2,dmsgs1))
            dattrs1 ]
      net(ddmsgs0,ddmsgs1)  conf1}
    if getTgt(msgd) == epid
    /\ toRcv(dmsgs) dattrs1 := decodeDialect(dattrs,msgd)
    /\ toSend(dmsgs2) devatts1 :=
        (if dmsgs :: DMsg
         then rcvMsg(epid, devatts, DMsg2Msg(dmsgs))
         else toSend(mtDM) devatts
         fi)
    /\ conf1 := doLog(conf,devatts1)
    /\ devatts2 := clearToLog(devatts1) } .
In addition to the dialect send and receive rules, the auxiliary functions mte and passtime used by the rule labeled tick are extended to account for the nesting of configurations. If there are messages in the local network then they need to be processed before time can pass. These messages will be messages from epid the should be dialected and placed in the global network.
  eq mte(atts conf([epid | devattrs] net(dmsgs0,dmsgs1)),ni)
   = (if dmsgs0 dmsgs1 =/= mtDM
      then 0
      else mte(atts, mte(devattrs,ni))
      fi) .
Similarly, the effect of passing time is simply propagated to the encapsulated agent, ignoring the local network element.
  eq passTime(conf([epid | devattrs] net(dmsgs0,dmsgs1)),nz)
  conf([epid | passTime(devattrs,nz,mtA)]  net(dmsgs0,dmsgs1)) .
7.3 CoAP dialect transform
As a step towards realizing the dialects as theory transformations, we define a function D that maps a term representing an initial system configuration to its dialected form. D calls the auxilliary function DX with the system term to transform and the set of ids of non-attacker endpoints computed by getIds. The function DX uses DAs to transform the individual agents of the configuration aconf, and adds the initial network element. The argument "xxxx" of DAs is used to construct abstract representations of the pairwise shared secrets. The last argument of DAs is a configuration accumulator. We also define a partial inverse, UD, to the dialect transform D. The partial inverse is only meaningful when applied to systems with no dialected messages pending. It is generally applied only to inital or terminal states, which have empty network elements.
    op D : Sys -> Sys .
    op DX : Sys Strings -> Sys .
    eq D(sys) = DX(sys,getIds(sys)) .
    ceq DX({aconf net(mtDM,mtDM)},eids) =
       { daconf net(mtDM,mtDM) }
    if daconf := DAs(aconf,eids,"xxxx",mt) .
The function DAs applies the local agent transform DA to each agent with identifier in the set eids and leaves any other agent (i.e. an attacker) unchanged. The function DA does the actual wrapping of a protocol role/endpoint producing a meta-agent with the same identifier and a conf attribute containing the original agent with its local network stubb. The attributes of the meta-agent wrapper consist of the attributes, sharedDialectAttrs that are the same for all wrappers, a set of seed attributes, a pair for each possible communication partner (network link) and a set of index attributes, one for each possible communication partner.
    op DA : Agent Strings String -> Agent .
    ceq DA([eid | devatts],eids,str) =
        [eid | conf([eid | devatts]
                    net(mtDM,mtDM)) ddevatts]
    if ddevatts := sharedDialectAttrs
                   mkSeeds(eid,eids,str,mtA)
                   mkIxCtrs(eids,mtA) .
To accomodate general scenarios, we assume any pair of non-attacker agents might communicate. When called from DX the str will be "xxxx" and result of mkSeeds will be a set of attribute pairs seedFr(eid,"xxxx" + eid1 + eid) and seedTo(eid,"xxxx" + eid + eid1) for each eid in eids. Similarly the function mkIxCtrs produces an attribute ixCtr(eid,0) for each eid in eids.
As an example, consisder the initial system
initSys =
 { net(mtDM, mtDM)
  ["dev0" | w4Ack(mtDM) w4Rsp(mtM) ...
            rsrcs(mtR) ctr(0)
            sendReqs(...)  config(...) sndCtr(0)]
  ["dev1" | w4Ack(mtDM) w4Rsp(mtM) ...
            rsrcs(rb("door", "unlocked")) ctr(0)
            sendReqs(nilAM) config(...) sndCtr(1)]
  ["eve" | kb(mtDM) caps(drop)]}
Its dialect transform, D(initSys) is
{ net(mtDM, mtDM)
  ["dev0" | conf(net(mtDM, mtDM)
          ["dev0" | w4Ack(mtDM) w4Rsp(mtM) ...
                    rsrcs(mtR) ctr(0)
                    sendReqs(...) config(...) sndCtr(0)])
    used(mtU) randSize(128)
        seedTo("dev1", "xxxxdev1dev0")
        seedFr("dev1", "xxxxdev0dev1") ixCtr("dev1", 0)]
  ["dev1" | conf(net(mtDM, mtDM)
            ["dev1" | w4Ack(mtDM) w4Rsp(mtM) ...
                     rsrcs(rb("door", "unlocked"))
                     ctr(0) sendReqs(nilAM) config(...) sndCtr(1)])
    used(mtU) randSize(128)
        seedTo("dev0", "xxxxdev0dev1")
        seedFr("dev0","xxxxdev1dev0") ixCtr("dev0", 0)]
  ["eve" | kb(mtDM) caps(drop)]}
The function UD extracts CoAP endpoints from conf attributes and passes other configuration elements unchanged
op UD : Sys -> Sys .
op UDX : Conf Conf -> Conf .
eq UD({conf}) = {UDX(conf,mt)} .
eq UDX([eid | conf([eid | devatts] conf1) ddevatts] conf0,
               uconf) =
      UDX(conf0, uconf [eid | devatts]) .
eq UDX(conf0, uconf) = uconf conf0 [owise] .
**** produce configurations by omitting the final system construction
op UDC : Conf -> Conf .
**** eq UDC(conf net(dmsgs0,dmsgs1)) = UDX(conf,mt) .
eq UDC(conf) = UDX(conf,mt) .
Definition 2 (Initial dialected CoAP configuration)
An initial dialected CoAP system, dsysI, is the result of transforming an initial CoAP system sysI
$\mathtt{dsysI}=\mathtt{D}(\mathtt{sysI}).$
7.4 Dialected CoAP Scenarios
As we prove in Appendix 0.C dialecting mitigates attacks of a reactive attacker. To illustrate this claim we can lift the scenarios for reactive attacks of Section refsec:reactive-attacks to dialected scenarios and lift the search for attacks to the dialected situation. To do this we apply the transform D to the initial system term and apply the inverse (on configurations) UDC to uses of the search target pattern c:Conf in the search condition. For example in the R1 attack (the attacker locks the door after the client has unlocked it) example the search
    search({raR1(5,0,10,false)}) =>! {c:Conf} such that
        checkRsrc(c:Conf,"dev1","door","lock") .
when lifted to a dialected system becomes
    search D({raR1(5,0,10,false)}) =>! {c:Conf} such that
        checkRsrc(UDC(c:Conf),"dev1","door","lock") .
There are 2 solutions in 101 states visited in the original search while no solution is found among 121 states visited in the search for the dialected system.
As one more example recall the example of R2 type attacks, where a client is invoking server proceses one at a time and the attacker manages to start some servers early, thus having more than one server active concurrently. The original search is
    search iSysX(3,0,caps-1) =>+ sys:Sys such that
          size(epswrb(sys:Sys,rb("sig","on"))) > 1 .
and its dialected form is
    search D(iSysX(3,0,caps-1)) =>+ sys:Sys such that
        size(epswrb(UD(sys:Sys),rb("sig","on"))) > 1 .
There are 132 solutions from 767 states visited in the undialected scenario while in the dialected case no solution is found in 553 visited.
The remaining cases are similar.
8 Dialect properties
Here we consider the relation of traces of an initial system $\mathtt{sysI}=\{\mathtt{confI}\}$ running the CoAP messaging protocol specification to traces of the corresponding dialected system D(sysI) in the presence of the different attack capabilities, including mtC (none). Although CoAP is designed to run over UDP (unreliable transport) we also consider the case of reliable transport. We note that the specified dialect wrapper (i) drops messages that fail decoding, and (ii) has a notion of used lingo parameter such that messages with previously used lingo parameters are dropped. We consider instances of the following relation:
$\mathsf{D}(\mathtt{mkISys}(\mathsf{att}(c)))~{}\sim?~{}\mathtt{mkISys}(\mathsf{att}(c))$
where $\mathtt{mkISys}(\mathsf{att}(c))$ denotes an initial system configuration with endpoints in their inital state, an empty network, and an attacker with capabilities $c$. The question is what is ?similar-toabsent?$\sim$?∼ ?. One candidate is stuttering bisimulation ($\approx$). As a second candidate we say that a dialect $\mathsf{D}$ is strongly attack reducing for attack messages M iff for any trace of $\mathtt{mkISys}(\mathsf{att}(c))$ if an attacker generated message M is first delivered at step $j$ then there is a corresponding trace of $\mathsf{D}(\mathtt{mkISys}(\mathsf{att}(c)))$ with M is delivered at the corresponding step $j^{\prime}$ and is dropped. In some more detail we define
$\mathsf{D}(\mathtt{mkISys}(\mathsf{att}(c)))\precsim[rcv[M]]~{}\mathtt{mkISys}(\mathsf{att}(c))$
to mean any trace from the lhs without the an attacker event has a corresponding trace from the rhs, and dually. If $\tau$ is a trace of either side with an event $\mathtt{rcv}[M]$ the first occurrence of receive of an attacker produced message, then the lhs trace drops the offending message and any corresponding rhs trace processes the message and possibly diverges from the lhs. In the cases considered this relation becomes
$\mathsf{D}(\mathtt{mkISys}(\mathsf{att}(c)))\approx\mathtt{mkISys}(\mathsf{att}(\mathtt{drop})).$
Dialect basic correctness.
In the absence of attacks in a reliable or unreliable network the dialect wrapper simply encodes messages sent by an endpoint, and decodes them before passing them to the receiver endpoint. Thus a dialected CoAP system is stuttering bisimilar to the original system.
$\mathsf{D}(\mathtt{mkISys}(\mathsf{att}(\mathtt{mtC})))\approx\mathtt{mkISys}(\mathsf{att}(\mathtt{mtC}))$
For proof see Appendix 0.C Theorem 0.C.1.
Dialecting over unreliable transport.
Here we consider the question: running over an unreliable transport, how do $\mathsf{D}(\mathtt{mkISys})$ in the presence of a given attack capability, $\mathtt{cap}$, and $\mathtt{mkISys}(\mathsf{att}(c))$ relate?
$\mathsf{D}(\mathtt{mkISys}(\mathsf{att}(c)))~{}\sim?~{}\mathtt{mkISys}(\mathsf{att}(mtC))$
The following is an informal analysis giving the intuitions. For proof see Appendix 0.C Theorem 0.C.2.
  • (i)
    $c=\mathtt{drop}$ or $c==\mathtt{delay}(n)$. In this case ?similar-toabsent?$\sim$?∼ ? is stuttering bisimilar since the network can drop or delay messages as well! What the attacker can do is significantly change the probability profile of drops/delays. [^3] This suggests a stronger relation than bisimilarity might be useful
    $\mathsf{D}(\mathtt{mkISys}(\mathsf{att}(c)))\approx\mathtt{mkISys}(\mathsf{att}(mtC))$
  • (ii)
    $c=\mathtt{divert}$ (edit source, or destination or both). Assuming each communicating pair in the dialected system has a unique shared secret, $\mathsf{D}(\mathtt{mkISys}(\mathsf{att}(c)))$ will drop the diverted message, thus making it look like a network drop. Without dialecting the new receiver might handle the diverted message, possibly causing problems, for example the temperature example (reading room vs oven temperature).
    $\mathsf{D}(\mathtt{mkISys}(\mathsf{att}(\mathtt{divert})))\approx\mathtt{mkISys}(\mathsf{att}(mtC))$
    Note that if all three endpoints involved share the same secret (the initial seed) then the dialect layer will not detect the redirection.
  • (iii)
    $c=\mathtt{replay}(n)$. There are two cases: (a) the original message wasn’t delivered—the replay appears as a delay or resend and won’t be dropped by the wrapper and will correspond to a trace on the rhs where the original message was delivered; (b) the original message was delivered—the replay is dropped. [^4] Note that if the lingo policy is to allow a parameter to be used 2 or more times before rejecting, then the second use of a parameter may be rejected after a replay causing dialecting to be complicit in the attack. Thus we chose a single-use policy.
    $\mathsf{D}(\mathtt{mkISys}(\mathsf{att}(\mathtt{replay}(n))))\approx\mathtt{mkISys}(\mathsf{att}(mtC))$
  • (iv)
    $c=\mathtt{edit}$ a message component (not implemented yet). There are two cases:
    • (iv.a)
      the message component is dialected, assuming the attacker can’t break the dialecting, undialecting will fail and the message will be dropped.
      $\mathsf{D}(\mathtt{mkISys}(\mathsf{att}(\mathtt{edit}(a))))\approx\mathtt{mkISys}\mathtt{mkISys}(\mathsf{att}(mtC))$
    • (iv.b)
      the message component is not dialected, then dialecting doesn’t change anything
      $\mathsf{D}(\mathtt{mkISys}(\mathsf{att}(\mathtt{edit}(b))))\approx\mathtt{mkISys}(\mathsf{att}(\mathtt{edit}(b)))$
  • (v)
    $c=\mathtt{create}$ (message creation not implemented yet). Assuming the attacker can’t break the dialecting, undialecting will fail and the message will be dropped
    $\mathsf{D}(\mathtt{mkISys}(\mathsf{att}(\mathtt{edit}(\mathtt{create}))))\approx\mathtt{mkISys}(\mathsf{att}(mtC))$
Dialecting over reliable transport.
Now we consider the question: In the presence of given attack capability, $c$, running over a reliable transport, how do $\mathsf{D}(\mathtt{mkISys}(\mathsf{att}(c)))$ and $\mathtt{mkISys}(\mathsf{att}(c))$ relate?
$\mathsf{D}(\mathtt{mkISys}(\mathsf{att}(c)))~{}\sim?~{}\mathtt{mkISys}(\mathsf{att}(c))$
  • (i)
    $c=\mathtt{drop}$ or $c==\mathtt{delay}(n)$. Dialecting doesn’t affect drop or delay attacks.
    $\mathsf{D}(\mathtt{mkISys}(\mathsf{att}(\mathtt{drop},\mathtt{delay}(n))))\approx\mathtt{mkISys}(\mathsf{att}(\mathtt{drop},\mathtt{delay}(n)))$
  • (ii)
    $c=\mathtt{divert}$ (edit source, or destination or both of message M. The dialect wrapper will drop a diverted message, as it is coded with the wrong secret. The base CoAP receiver may process the message.
    $\mathsf{D}(\mathtt{mkISys}(\mathsf{att}(\mathtt{divert})))~{}\precsim[rcv(M)]~{}\mathtt{mkISys}(\mathsf{att}(\mathtt{divert}))$
  • (iii)
    $c=\mathtt{replay}(n)$ of M. Note that with reliable transport, the original message will have been delivered. The dialect wrapper will drop the message, as its lingo (parameter) is already used.
    $\mathsf{D}(\mathtt{mkISys}(\mathsf{att}(\mathtt{replay}(n))))\approx\mathtt{mkISys}(\mathsf{att}(\mathtt{mtC}))$
    • (iii.a)
      $n$ is small enough that $M$ is delivered within the original message’s (identifier) lifetime , the CoAP receiver may repeat the original response, but will not process any message requested actions.
      $\mathsf{D}(\mathtt{mkISys}(\mathsf{att}(\mathtt{replay}(na))))\approx\mathtt{mkISys}(\mathsf{att}(\mathtt{replay}(na)))$
    • (iii.b)
      $M$ is delivered after the original message’s lifetime expires. In this case the CoAP receiver will process the replay.
      $\mathsf{D}(\mathtt{mkISys}(\mathsf{att}(\mathtt{replay}(nb))))\precsim[\mathtt{rcv}(M)]~{}\mathtt{mkISys}(\mathsf{att}(\mathtt{replay}(nb)))$
  • (iv)
    $c=\mathtt{edit}$ a component if message M (not implemented yet). event = rcv(M), There are two cases:
    • (iv.a)
      the edited message component is dialected. Assuming the attacker can’t break the dialecting, undialecting will fail and the message will be dropped, while the undialected system will attempt to process the message.
      $\mathsf{D}(\mathtt{mkISys}(\mathsf{att}(\mathtt{edit}(a))))\precsim[rcv[M]]~{}\mathtt{mkISys}(\mathsf{att}(\mathtt{edit}(a)))$
    • (iv.b)
      the message component edited is not dialected, then dialecting doesn’t change anything.
      $\mathsf{D}(\mathtt{mkISys}(\mathsf{att}(\mathsf{att}(\mathtt{edit}(b)))))\approx\mathtt{mkISys}(\mathsf{att}(\mathtt{edit}(b)))$
  • (v)
    $c=\mathtt{create}$ message M. Assuming the attacker can’t break the dialecting, the dialect wrapper will drop the message, but $\mathtt{mkISys}(\mathsf{att}(\mathtt{create}))$ will attempt to process it.
    $\mathsf{D}(\mathtt{mkISys}(\mathsf{att}(\mathsf{att}(\mathtt{create}))))\precsim[rcv[M]]~{}\mathtt{mkISys}(\mathsf{att}(\mathtt{create}))$
9 Related Work
The work most closely related to the present work is [GEMM-2023esorics], which we will refer to as ESORICS23 in the following. Complimentary work on dialects is discussed briefly in Section 2. We refer to ESORICS23 for a more complete review of previous work on dialects.
ESORICS23 addresses three key aspects of dialects: synchronization mechanisms; protocol and lingo genericity; and attack model vs dialect choice. It also defines taxonomies for key dimensions of a dialect.
A general framework is proposed defining notions of dialect and lingo as (parameterized) transformations on protocol theories, where a protocol theory is a generalized actor theory with a sort Configuration and a constant initConf of sort Configuration, representing the initial protocol system state. A lingo defines a pair $f,g$ of parameterically inverse functions for obfuscating and deobfuscating protocol messages: $g(f(m,p),p)=m$. A lingo transforms a protocol theory to a new protocol theory by adding the function definitions. The dialect transform operates on the protocol theory together with the lingo transformed theories. In the resulting theory, protocol objects are wrapped with dialect meta objects with the same identifier and rules are provided for sending/receiving messages at the dialect level.
The CoAP Dialect presented in Section 7 has one lingo, defined in the module COAP-DIALECT-LINGOS which extends the functional part of the CoAP specification. The CoAP dialect is specified as the result of the transform, by including the lingo and protocol modules and adding rules for receiving and sending messages. The initial configurations are defined in protocol and dialect test modules, where dialected configurations are obtained as in the ESORICS23 case by wrapping protocol objects by meta objects.
We note that the definition of protocol theory with a fixed initial configuration doesn’t quite work if you want it to support a variety of scenarios. For example MQTT and CoAP support networks of almost arbitrary size and topology. One can define a protocol theory for each scenario of interest, with a dialect being generic with respect to the initial configuration. But this looses the information that all the scenarios use the same rules and data types thus seems not quite satisfactory. Perhaps a notion of scenario constructor parameterized by a network topology and a means of specifying initial shared secrets could be useful. Also, configurations are ‘open’ in the sense that each actor or subset of actors and messages is a configuration and a configuration is a subconfiguration of infinitely many larger configurations. In most cases rewrite rules apply to a subconfiguration, for example an actor and a message, independent of what configuration it is part of. However, some rules need to know that they are seeing the full configuration. This can be achieved by introducing the notion of system as an encapsulated configuration. The question is whether the notion of system should be a first class concept for protocols and dialect transformations.
Synchronization mechanisms.
Synchronization can happen on different time scales: when the set of lingos in use changes, and when current lingo parameters change. The latter is likely to be more frequent than the former, although these are not necessarily distinct, since a lingo parameter could be use to switch between lingo functions say xor versus shuffle.
Two classes of synchronization mechanism are identified in ESORICS23: periodic–that uses synchronized clocks and fixed time intervals for changing to the next phase and its lingo set) or next parameter values; and aperiodic–using information in the local history to determine when/what to change.
In the CoAP dialect studied here, there is one (abstract) lingo with a changing parameter that selects the next pseudo random string in a practically unpredictable (i.e. hard to predict) sequence. The CoAP dialect uses aperiodic synchronization for reasons discussed in Section 6. The relevant history is a set of counters one for each communication partner that is incremented upon each use.
Genericity.
In the framework proposed in ESORICS23, dialects are parametric in a list of lingos and thus can be applied to any protocol theory for which the lingos are applicable. Assuming that message content is of type string enables a wide choice of lingos independent of the meaning of the strings. The CoAP lingo functions assume an abstract message content type Content and thus the CoAP dialect could be applied to any protocol specification with this content type (and suitable configuration sort). We note that the example used in ESORICS23 suggests that the behavior of the dialect depends on the message type and even uses information from the CONNECT messages, while the CoAP dialect does not look inside message content. Presumably some dependence on message structure could be made a parameter to the dialect in addition to the lingo parameters.
Attack model.
The ESORICS23 attack model starts with two models proposed in [ren-etal-2023seccom]: the off-path attacker that can send and receive messages, but can not see messages exchanged between its targets; and the on-path attacker with all the capabilities of the off-path attacker, plus the ability to observe communications between honest endpoints. ESORICS23 adds a third model: active on-path attacker with the capabilities of the on-path attacker plus the ability to block, redirect, and alter message content. A refinement is also introduced–(m, n, t) attacker that in time t can observe m messages and block n messages (and send/receive, block, …). The goal of a dialect is low probability that an “attacked” message will be accepted by intended receiver.
In this paper we define two attack models for CoAP messaging systems: active and reactive. In either case, an attacker is allowed a bounded set of actions from a set of capabilities (depending on the model). The bound on number of actions is an alternative to the bounds in the (m,n,t) attack model as a means to limit the attacker strength. The active attacker is based on vulnerabilities identified in [coap-attacks]. The capabilities are: block/drop, delay, replay, or redirect messages. Our active attack model differs from the active on-path model in that attackers don’t receive or create messages. Like [coap-attacks] we assume communications may be encrypted, so focuses on attacks that don’t involve parsing messages – all the attacker learns from a message is the target, source and a content blob.
The ESORICS23 paper is focused on understanding the structure of dialects and attack models and gives some examples and informal statements about the security guarantees of lingos and dialects. In the present paper we formally specify an attack model and carry out reachability analyses concerning specific undesired states of simple applications using the CoAP protocol (inspired by [coap-attacks]). We also prove properties relating executions of dialected and undialected scenarios with and without attacks.
Taxonomies.
ESORICS23 proposes a taxonomy of dialects with four classes: Protocol Specific; Protocol Generic; Protocol Modifying; Protocol Wrapping. The CoAP dialect is essentially generic for single lingo case (the lingo needs to be suitable for the protocol message structure) and it is Wrapping.
A lingo taxonomy is proposed with classes Shuffle/permutation (assumes messages are bit strings); Split; Flow (via customized messages); Crypto (HMAC, XOR, …); and Use of unused/random field. The CoAP lingo is abstract, the functions could be realized using (combinations of) shuffle, xor, or crypto. We did not consider Splitting as one of the aims of CoAP is to avoid packet fragmentation. With an unreliable transport, the dialect/lingo itself would have to do some extra work to avoid additional dropped messages.
Other points.
It appears that in the ESORICS23 framework all honest actors (protocol actors) share the same secrets. Examples only consider two actor systems, so its not completely clear. In section 8 we show examples of systems with 3 or more endpoints where its important to have distinct pairwise shared secrets.
ESORICS23 also proposed notions of (generic) vertical and horizontal composition. That is beyond the scope of the present work, but certainly an interesting future direction along with partial evaluation transformations to eliminate extra overhead where resources are very limited.
Complimentary to ESORICS23, the current paper carried out the dialect theory transformation by hand, formalizing only the wrapping (and unwrapping) of system configurations, and focused effort on formalizing attack models and carrying out formal reachability analyses for a variety of scenarios.
10 Conclusions
Dialects such as ESORIC23 or the CoAP dialect can protect against replay, editing message source and/or target, and message forging. They can not protect against interference attacks such as dropping or delaying. These dialects also don’t protect against piggy back attack. In this attack there are two attackers X,Y where X wants to send signals to Y. When A sends M to B, X intercepts and sends to Y (on port j). Y learns the signal represented by j and forwards M to B.
This suggested the reactive attacker model for dialects–a non-interference model. The attacker can observe (copy), construct/edit, transmit (no block/drop, delay) thus can replay, redirect, spoof. We showed that dialects satisfying minimal conditions provide security against non-interference attacks in the sense that a dialected protocol in the presence of such attacks is bisimilar to the protocol alone.
We conclude with so ideas for future work. It could be interesting to consider dialects that emit apparently random messages between themselves? These could be versions of previously sent messages or made up messages of suitable size. They should be similar to ordinary traffic and not sent too often. Can this obfuscate patterns the attacker relies on? Another alternative is deploying intermediate nodes with a dialect layer only. These would be used to obfuscate where the message is actually going. There is still work to be done to capture a generic attack model for messaging protocols: what are the properties that applications care about? how to represent them generally? what attack capabilities are needed to succeed in violating critical properties? is there a notion of complexity of property that corresponds to a minimal bound on successful attacker capability?
Acknowledgments.
Talcott thanks Dr. Catherine Meadows for many insightful discussions that lead to some of the attack model ideas and other results. This material is based upon work supported by the Naval Research Laboratory under Contract No. N00173-23-C-2001. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the Naval Research Laboratory.
Appendix 0.A Characterization of reachable CoAP system configurations
As for real-time Maude specifications [olveczky08tacas] the rewrite rules for the CoAP specification (and it dialected form) are either instantaneous rules or the tick rule that models passing of time. The minimum time to elapse function, mte(sys), determines when time can pass. The following propositions give essential properties of reachable CoAP and dialected CoAP system configurations according to the value mte(sys).
Proposition 1 (Mte based characterization of reachable CoAP configurations)
Let sysI be an initial configuration and let sysC be a CoAP system configuration reachable from sysI. Then mte(sysC) is either 0, nz a non-zero natural number, or the constant infty (representing infinity) and one of the following holds.
  • 1.
    mte(sysC) = nz:NzNat. In this case the tick rule is enabled (and no other rule is enabled (except possibly net which commutes with and is independent of tick). Further more
    • for all msg @ d in the network, $\mathtt{d}>0$ (ow mte is 0).
    • for all msg @ d in some w4Ack attribute $\mathtt{d}>0$ (ow mte is 0).
    • if the sendReq attribute of some endpoint is not empty (nilAM) then either there are more than the w4Ack limit of requests waiting for acknowledgement (and hence rule devsnd is not enabled), or the corresponding sndCtr attribute is non-zero.
  • 2.
    mte(sysC) = 0 In this case some rule is enabled according to one of the following cases.
    • there is a message with zero delay in the net, then the rule rcv is enabled (possibly preceeded by net) to move the message from network input to network output;
    • there is a w4Ack attribute of some endpoint with a zero delay message, then the ackTimeout rule is enabled;
    • there is a sendReqs attribute of some endpoint that is not empty and sending is enabled, so the rule devsnd is enabled.
  • 3.
    mte(sysC) = infty. In this case sysC is a terminal configuration, no rules apply since the net is empty, all w4Ack attributes are empty, and all sendReqs are empty buy definition of mte.
The definition of mte is extended to dialected CoAP configurations by adding the rule that mte is 0 if any local nets are non-empty. Thus messages in local nets must be delivered or dialected and put in the global net before time can pass. The consequences are captured in the next proposition.
Proposition 2 (Mte based characterization of reachable dialected CoAP configurations.)
As for the non-dialected case there are three cases.
  • 1.
    $\mathtt{mte(dsysC)}=\mathtt{nz:NzNat}$. In this case the value of mte is determined by the global network and the sysC level attributes and only the tick rule (and possibly net) is enabled.
  • 2.
    $\mathtt{mte(dsysC)}=\mathtt{0}$. In this case, if there is a message some local net the either ddevsnd or rcv is enabled. Otherwise either ddevrcv is enabled or some rule at the sysC level as shown in Proposition 1.
  • 3.
    $\mathtt{mte(dsysC)}=\mathtt{infty}$. In this case dsysC is terminal. As for sysC the net is empty (global and local), w4Ack attributes are empty, and sendReq attribues are empty so no rule is enabled.
Appendix 0.B CoAP vulnerabilities
The paper [coap-attacks] motivates an extension to the CoAP standard by describing several vulnerabilities where attacks do not depend on being able to see message content, i.e. these attacks can in principle be carried out against CoAP running over DTLS or other transport providing encryption.
Four classes of vulnerability are discussed using message diagrams to illustrate attack scenarios. Three of the classes rely only on the ability to drop or delay messages. In some cases the vulnerability is due to reuse of tokens resulting in requests connected to multiple responses. These scenarios do not exhibit vulnerabilities in the Maude specification, but in some cases we give alternative scenarios that realize a similar vulnerability. With the execption of the example that relys on mismatching responses to GETs of different resources on one server, dialecting does not change the success of attacks as it can not detect blocking/dropping or delay. Our modified version of the mismatch case uses two servers and in this case dialecting prevents the attack becuase lingos are pairwise disjoint. The fourth class concerns message fragmentation in the network. This is not modeled in the current Maude CoAP specification.
In the following we show how the three relevant vulnerabilities can be formally modeled and analyzed using the Maude specification. In the diagrams, the vertical lines are participant send/receive event timelines. The client is labelled dev0, the server labelled dev1 and the attacker labelled eve. The state of the server resource "door" is indicated at the beginning and end of the server timeline. In a timeline, os mark normal send or receive events, X indicates a message drop, and @ indicates a message capture (for later release).
We use the function tCS2C defined in Section 3.4 to construct scenario initial system configurations. For convenience we repeat the definition here.
op tCS2C : Nat Nat AMsgL RMap Nat Nat
           Nat Nat AMsgL RMap Nat Nat
          Caps -> Conf .
eq tCS2C(n0,j0,amsgl0,rbnds0,mqd0:Nat,w4ab0:Nat,
         n1,j1,amsgl1,rbnds1,mqd1:Nat,w4ab1:Nat,
        caps) =
   net(mtDM,mtDM)
   mkDevC(n0,j0,amsgl0,rbnds0,mqd0:Nat,w4ab0:Nat)
   mkDevC(n1,j1,amsgl1,rbnds1,mqd1:Nat,w4ab1:Nat)
   (if caps == mtC then mt else mkAtt(caps) fi)  .
It generates a configuration with two endpoint agents (constructed using mkDevC) and an attacker (if caps is non empty). The arguments specify initial application messages (amsglj) and resource maps (rbndsj). The first two numbers for a device construction determine the device id and inital delay for message sending. The last two numbers for a device construction specify the congestion control behavior: a new message can not be sent if there are more than w4ab requests not yet acknowledged, and there must be a delay of a least mqd between requests.
We also use three application message constructors to generate requests. [^5] Recall that a client will retry a confirmable request if no acknowlegment or response is received within a specified time. Non-confirmable requests are not acknowledged and are not retried.
  • mkPutN(tag,server,resource,val): a non-confirmable request to server to set resource to val.
  • mkPutC(tag,server,resource,val): a confirmable request to server to set resource to val.
  • mkGetN(tag,server,resource) a non-confirmable request to get the resource value from server.
See Section 3.4 for definitions.
To check if an attack is possible in a given scenario, we search from the constructed initial configuration initSys to a terminal configuration  [^6] See section 0.A for a characterization of terminal configurations. that satisfies a property P(c:Conf, params) characterizing the effects of the attack typically in terms of properties of unperturbed executions that are violated. Requiring solutions to be terminal allows us to rely on a number of consequences such as if a message is sent, then it will be received unless it is dropped.
search initSys =>! {c:Conf ["eve" | a:Attrs caps(mtC)]}
    such that P(c:Conf, params)
This lists all the ways the property P can be satisfied. In some cases, we also search for executions in which the attacker fails, i.e. the expected execution properties hold. Including the term ["eve" | a:Attrs caps(mtC)] in the search pattern ensures that the attacker will use all its capabilities in reaching a solution state.
We use the following predicates to form the search properties. These are very basic properties that characterize the state of endpoints at the end of a scenario.
Properties of the server state.
  • checkRsrc(c:Conf,sv,r,v): The resource r of server sv has value v.
  • hasRspTSnt(c:Conf,sv,cl,aid): Server sv has sent a response to client cl in response to a request with token matching aid (the application provided message identifier, which is part of the message id and token strings).
  • rspTSntBefore(c:Conf,sv,cl,aid0,aid1): Server sv sent responses to client cl in response to a request with token matching aid0 and a requesst with token matching aid1 with the aid0 matching response being sent first.
Properties of the client state.
  • hasRspTRcd(c:Conf,cl,sv,aid): Client cl had received a response from server sv to a request with token matching aid. (The empty string matches any token.)
  • hasGetRsp(c:Conf,cl,sv,aid,val): Client cl had received a response from server sv to a GET request with token matching aid. The response body contains value val. (If val is the empty string, the value check is omitted.)
  • rspPend(c:Conf,cl,sv,aid): Client cl is waiting for a response from server sv to a request with token matching aid.
The following subsections treat the different vulnerabilities. They begin with a diagram of the proposed attack, followed by a formal definition of the corresponding scenario, search commands and results demonstrating attack success and failure situations. We show the scenario definitions and search commands for the original CoAP case. The search command for the corresponding dialected scenario is given by applying the dialect transformation D (defined in Section 7.3) to the initial system configuration and lifing the search pattern to the dialected structure to expose the relevant elements. Thus the dialected version of the CoAP search command
        search initSys =>!
          {c:Conf ["eve" | a:Attrs caps(mtC)]
                  [epid | devatts:Attrs]}
          such that prop(c:Conf,epid) .
is
       search D(initSys) =>!
         {c0:Conf  ["eve" | a:Attrs caps(mtC)]
                   [epid | conf(c:Conf)]
          ddattrs0:Attrs]}
        such that prop(c:Conf,epid) .
Note that the goal property doesn’t change.
0.B.1 The Selective Blocking Attack.
This scenario illustrates the ability of an attacker to block a single message. How the attacker recognizes the intended target is not discussed in [coap-attacks]. Th scenario is adapted from Figures 1,2 of [coap-attacks].
      dev0      eve    dev1        dev       eve     dev1
      ---      ---   unlock        ---       ---    unlock
       o -PUTN->X      |            o -PUTN->  ------>  o
                       |            |         X <-2.04- o
                    unlock          |                 lock
          (1) drop message            (2) drop response
          Figures 1,2 scenarios
-PUTN-> a non-confirmable request to lock the door. -2.04-> the PUTN response. (1) Attacker blocks the request. (2) Attacker blocks the response.
The function caFig1.2 specifies the scenario for the message dropping attacks of Figures 1 and 2, parameterized by congestion control parameters (mqd:Nat the minimum time between message sends; more that w4b:Nat requests in w4Ack blocks sending additional requests).
    op caFig1.2 : Nat Nat -> Conf .
    eq caFig1.2(mqd:Nat,w4b:Nat) =
         tCS2C(0,0,mkPutN("putN","dev1","door","lock"),
                   mtR,mqd:Nat,w4b:Nat,
               1,1,nilAM,rb("door","unlocked"),2,0,
               drop) .
Figure 1 successful attack.
Dropping a request is specified by checking that the server did not make any response (hence it did not receive the request) and the client awaits a response (it sent a request). Note this is redundant given the semantics, but a systematic characterization.
    search {caFig1.2(5,0)} =>!
      {c:Conf ["eve" | a:Attrs caps(mtC)]} such that
      not(hasRspTSnt(c:Conf,"dev1","dev0","putN")) and
      rspPend(c:Conf,"dev0","dev1","putN")  .
or by checking that the server resource has the original value.
      search {caFig1.2(5,0)} =>!
      {c:Conf ["eve" | a:Attrs caps(mtC)]} such that
      checkRsrc(c:Conf,"dev1","door","unlocked") .
In either case there are 2 solutions among 7 states. The results for the corresponding two dialected cases are similar, 2 solutions among 9 states visited.
Figure 2 successful attack.
Dropping the response can be specified by checking that the server sent a response and the client did not receive one.
    search {caFig1.2(5,0)} =>!
    {c:Conf  ["eve" | a:Attrs caps(mtC)]} such that
      hasRspTSnt(c:Conf,"dev1","dev0","putN") and
      checkRsrc(c:Conf,"dev1","door","lock") and
      rspPend(c:Conf,"dev0","dev1","putN") .
There are 2 solutions out of 13 states visited.
Figure 1,2 attack failures.
To check that there are good executions in the presence of an attacker we check that the server sends a response and the client receives it and the attacker used all its capabilities.
    search {caFig1.2(5,0)} =>!
      {c:Conf ["eve" | a:Attrs caps(mtC)]} such that
         hasRspTSnt(c:Conf,"dev1","dev0","putN") and
         hasRspTRcd(c:Conf,"dev0","dev1","putN") .
There is no solution, some message has to be dropped and there are only two messages.
0.B.2 The Request Delay Attack.
This scenario is based on figure 3 of [coap-attacks].
          dev0        eve        dev1
          ----        ---        lock
           o -PUTNDU-> @
           o -PUTNSO->  ------>    o
           o <-------  <-2.01--    o
                       o -PUTNDU-> o
                       X <-2.04-   o
                                unlock
           Figure 3 scenario
The client intends to unlock the door, PUTNDU, then turn a signal on PUTSO. However the attacker delays the door unlock request and drops the PUTNDU response. The signal goes on with the door locked.
The Figure 3 attack scenario is constructed by the function caFig3.
      op caFig3 : Nat Nat Nat -> Conf .
      eq caFig3(d:Nat,mqd:Nat,w4b:Nat) =
       tCS2C(0,0,mkPutN("putND","dev1","door","unlock") ;
                 mkPutN("putNS","dev1","signal","on"),
                mtR,mqd:Nat,w4b:Nat,
            1,1,nilAM,rb("door","lock"),2,0,
            drop delay(d:Nat)) .
Figure 3 successful attacks.
From the diagram we characterize the attack by server properties saying a request with token matching putNS was received and processed before the request with token matching putND; the door is unlocked, and the client has not received a response to the message with token matching putND.
      search {caFig3(15,5,0)} =>!
      {c:Conf ["eve" | a:Attrs caps(mtC)]} such that
      checkRsrc(c:Conf,"dev1","door","unlock") and
      rspTSntBefore(c:Conf,"dev1","dev0","putNS","putND") and
      rspPend(c:Conf,"dev0","dev1","putND") .
There are 4 solutions in 330 states visited. In the dialected version of the scenario, there are 4 solutions among 496 states visited.
Figure 3 failed attacks.
Failed attacks are those executions where the attacker uses all its capabiities, but the endpoint events are as expected in absence of attacker. That is, the server has responded to putND before putNS and the client has received the two expected responses. If we require the client receive all responses there will be no solution, as something needs to be dropped. So we only require the client receive a response to putND.
      search {caFig3(15,5,0)} =>!
      {c:Conf  ["eve" | a:Attrs caps(mtC)]} such that
      checkRsrc(c:Conf,"dev1","door","unlock") and
      rspTSntBefore(c:Conf,"dev1","dev0","putND","putNS") and
      hasRspTRcd(c:Conf,"dev0","dev1","putND")  .
There are 4 solutions, for example the attacker might delay the putNS request, and drop the response to that request.
0.B.3 The Delay with Reordering Attack.
This scenario is based on Figure 4 of [coap-attacks]. Here the door is initially locked. The client requests unlock, possibly carries out some other interactions, and then requests lock. The attacker captures the unlock request. Since the unlock request is confirmable, the client resends it and the server sends a response to the resent request. The following lock request is processed as expected. Then the attacker releases the captured unlock request and drops the response. Thus the client believes the door is locked when in fact it is unlocked.
        dev0        eve        dev1
        ----        ---        lock
         o -PUTCU->  @
         o -PUTCU->  -------->  o  --- retry
         o <----     <---2.04-- o
         o -PUTNL->    ------>  o
         o <-------  <--2.04--  o
                     o -PUTCU-> o  --- release
                     X <-2.04-  o
                              unlock
           Figure 4 scenario
This scenario is constructed by the function caFig4x giving the attack a drop and a delay capability. The function is parameterized by the delay ammount (n:Nat) and the usual congestion controls.
      op caFig4x : Nat Nat Nat -> Conf .
      eq caFig4x(n:Nat,mqd:Nat,w4b:Nat) =
         tCS2C(0,0,mkPutC("putC","dev1","door","unlock") ;
                  mkPutN("putN","dev1","door", "lock"),
                  mtR,mqd:Nat,w4b:Nat,
             1,1,nilAM,rb("door","lock"),2,0,
             drop delay(n:Nat)) .
Figure 4 Successful Attack
The attack is specified by requiring the server to send a response to a request with token matching "putC" followed by a "putN" followed by another "putC" with the door unlocked at the end. We require the client to have received responses to "putC" and "putN" requests thus leading the client to believe it succeeded.
      search {caFig4x(10,5,0)} =>!
      {c:Conf ["eve" | a:Attrs caps(mtC)]} such that
      checkRsrc(c:Conf,"dev1","door","unlock") and
      hasRspTSnt(c:Conf,"dev1","dev0","putC") and
      hasRspTSnt(c:Conf,"dev1","dev0","putN") and
      rspTSntBefore(c:Conf,"dev1","dev0","putC","putN") and
      rspTSntBefore(c:Conf,"dev1","dev0","putN","putC") and
      hasRspTRcd(c:Conf,"dev0","dev1","putN") and
      hasRspTRcd(c:Conf,"dev0","dev1","putC") .
There are 8 solutions from 2600 states visited. In the dialected version there are 16 solutions from 10167 states visited.
Figure 4 Failed Attack.
The attacker can fail even though both capabilities are used. In this case we require the server response sequence to be simply putC then putN and the final door state to be "lock" and requier the client state as for the successful attack.
          search {caFig4x(10,5,0)} =>!
          {c:Conf ["eve" | a:Attrs caps(mtC)]} such that
          checkRsrc(c:Conf,"dev1","door","lock") and
          hasRspTSnt(c:Conf,"dev1","dev0","putC") and
          hasRspTSnt(c:Conf,"dev1","dev0","putN") and
          rspTSntBefore(c:Conf,"dev1","dev0","putC","putN") and
          hasRspTRcd(c:Conf,"dev0","dev1","putN") and
          hasRspTRcd(c:Conf,"dev0","dev1","putC") .
There are 45 solutions from 2586 states visited.
0.B.4 The Response Delay and Mismatch Attack.
This scenario is based on Figure 5 of [coap-attacks]. In this attack the client sends an unlock request, the attacker captures the response, since the resquest is non-confirmable the client doesn’t worry about the lack of response. Later the client sends a lock request that is dropped by the attacker, who releases the unlock response to trick the client into thinking the lock request succeeded, when in fact the door remains unlocked.
This works for clients that reuse tokens. It fails using the Maude client specification.
            dev0      eve       dev1
            ----      ---       lock
             o -PUTNU-->   -----> o
             |         @  <-2.04- o
             o -PUTNL->   X       |
             o <-2.04-    o       |
                               unlock
           Figure 5 scenario
This attack scenario is specified by the function caFig5x.
        op caFig5x : Nat Nat Nat -> Conf .
        eq caFig5x(d:Nat,mqd:Nat,w4b:Nat) =
           tCS2C(0,0,mkPutN("putNU","dev1","door","unlock") ;
                     mkPutN("putNL","dev1","door", "lock"),
                     mtR,mqd:Nat,w4b:Nat,
                 1,1,nilAM,rb("door","lock"),2,0,
                 drop delay(d:Nat)) .
Figure 5 intented attack fails
            search {caFig5x(10,5,0)} =>!
            {c:Conf ["eve" | a:Attrs caps(mtC)]} such that
            hasRspTSnt(c:Conf,"dev1","dev0","putNU") and
            not(hasRspTSnt(c:Conf,"dev1","dev0","putNL")) and
            checkRsrc(c:Conf,"dev1","door","unlock") and
            hasRspTRcd(c:Conf,"dev0","dev1","putNL") .
There is no solution. Also no solution in the dialected case.
Figure 5 alternative attack
The result that the client believes the door is locked when it is not can be realized by an attack where the unlock request is delayed and the unlock response is dropped.
          search {caFig5x(10,5,0)} =>!
          {c:Conf ["eve" | a:Attrs caps(mtC)]} such that
          hasRspTSnt(c:Conf,"dev1","dev0","putNU") and
          hasRspTSnt(c:Conf,"dev1","dev0","putNL") and
          rspTSntBefore(c:Conf,"dev1","dev0","putNL","putNU") and
          checkRsrc(c:Conf,"dev1","door","unlock") and
          hasRspTRcd(c:Conf,"dev0","dev1","putNL") .
There are 4 solution 4 among 330 states visited. In the dialected case there are 4 solutions among 499 states visited.
0.B.5 Delaying and mismatching response to GET requests
This attack is based on Figure 6 of [coap-attacks]. In this attack the attacker captures the response to a GET request, drops a subsequent GET request and releases the original response. Between GET requests the door state may have changed. A vulnerable client will accept the delayed response to the first GET as response to the second GET, and thus in the case where the door is locked initially, and becomes unlocked, client incorrectly thinks the door is locked, assuming the unlock request failed to arrive. The Maude client will accept the delayed response, but as response to the first GET and will not have any knowledge about the later state.
            dev0      eve        dev1
            ----      ---        lock
             o -GETN->   ------>  o    getN0
             |         @  <-lock- o
             o -PUTNU-->  ----->  |    putNU
             |         X  <-2.04- o
             o -GETN-> X          |    getN1
             o <-lock- o          |
                               unlock
            Figure 6 scenario
This attack scenario is constructed by the function caFig6x.
        op caFig6x : Nat Nat Nat -> Conf .
        eq caFig6x(d:Nat,mqd:Nat,w4b:Nat) =
           tCS2C(0,0,mkGetN("getN0","dev1","door") ;
                     mkPutN("putNU","dev1","door","unlock") ;
                     mkGetN("getN1","dev1","door") ,
                     mtR,mqd:Nat,w4b:Nat,
                 1,1,nilAM,rb("door","lock"),2,0,
                 drop drop delay(d:Nat)) .
Figure 6 intended attack fails.
To specify the intended attack we require that the server sends a response to "getN0" before sending a response to "putNU", at the end the door state is "unlock", and the client has a response to "getN1" with value "lock".
              search {caFig6x(10,5,0)} =>!
              {c:Conf ["eve" | a:Attrs caps(mtC)]} such that
              hasRspTSnt(c:Conf,"dev1","dev0","getN0") and
              hasRspTSnt(c:Conf,"dev1","dev0","putNU") and
              rspTSntBefore(c:Conf,"dev1","dev0","getN0","putNU") and
              checkRsrc(c:Conf,"dev1","door","unlock") and
              not(hasRspTSnt(c:Conf,"dev1","dev0","getN1") ) and
              hasGetRsp(c:Conf,"dev0","dev1","getN1","lock") .
There are no solutions, also no solutions in the dialected case.
Figure 6 alternate attack succeeds
In the alternate attack, the server requirement is the same as for the intended attack, but the client has a response to "getN0" but no response to either "putNU" or "getN1", thus the door is unlocked at the end, but the client does not know the state.
            search {caFig6x(10,5,0)} =>!
            {c:Conf ["eve" | a:Attrs caps(mtC)]} such that
            hasRspTSnt(c:Conf,"dev1","dev0","getN0") and
            hasRspTSnt(c:Conf,"dev1","dev0","putNU") and
            rspTSntBefore(c:Conf,"dev1","dev0","getN0","putNU") and
            checkRsrc(c:Conf,"dev1","door","unlock") and
            not(hasRspTSnt(c:Conf,"dev1","dev0","getN1") ) and
            hasGetRsp(c:Conf,"dev0","dev1","getN0","lock") and
            rspPend(c:Conf,"dev0","dev1","getN1") and
            rspPend(c:Conf,"dev0","dev1","putNU") .
There are 18 solutions in 2742 states visited. In the dialected case there are 18 solutions among 4675 states visited.
0.B.6 Delaying and mismatching response from other resource
This attack is based on Figure 7 of [coap-attacks]. In the original attack the client sends temperature requests to two resources (oven,room), and the attacker switches responses, dropping the the room temperature response, making the client think the room is on fire. We adapt this to the case of doors to two rooms, to keep to the door theme.
              dev0      eve        dev1
              ---       ---     d1/lock d2/unlock
               o  ----GETN(d1)--->   o  getN0
               |             @ <-d1- o
               o -GETN(p2)-> X       |  getN1
               o <-d1-       o       |
              Figure 7 scenario
With a vulnerable client the attack succeeds because the client uses same token for both GETs and will accept the response to the first GET as a response to the second GET request. The Maude client does not reuse tokens and this attack fails.
The attack scenario is constructed by the function caFig7x.
          op caFig7x : Nat Nat Nat -> Conf .
          eq caFig7x(d:Nat,mqd:Nat,w4b:Nat) =
             tCS2C(0,0,mkGetN("getN0","dev1","door1") ;
                       mkGetN("getN1","dev1","door2") ,
                       mtR,mqd:Nat,w4b:Nat,
                   1,1,nilAM,
                   rb("door1","lock") rb("door2","unlock"),2,0,
                   drop delay(d:Nat)) .
Figure 7 attack fails.
          search {caFig7x(10,5,0)} =>!
            {c:Conf ["eve" | a:Attrs caps(mtC)]} such that
          hasRspTSnt(c:Conf,"dev1","dev0","getN0") and
          not(hasRspTSnt(c:Conf,"dev1","dev0","getN1")) and
          hasGetRsp(c:Conf,"dev0","dev1","getN1","lock") .
There is no solution. Also no solution in the dialected case.
0.B.7 Delaying and mismatching response from other resource variant
This attack is a variant of the attack based on Figure 7 of [coap-attacks] using two servers rather than one server with two resources.
                dev0         eve      dev1    dev2
                ----         ---     unlock   lock
                o -GETN(1)--> #        |        |
                |             >  ---GETN(2)-->  o
                |             #  <---lock---    o
                o <-1:lock-   <        |        |
The client asks for the door state at dev1. The attacker redirects the request to dev2 and (un)redirects the response to appear to come from dev1. This scenario is generated by the function caFig7mod.
            op caFig7mod : Nat  Nat -> Conf .
            eq caFig7mod(mqd:Nat,w4b:Nat) =
                 tCS3C(0,0,mkGetN("getN0","dev1","door"),
                           mtR,mqd:Nat,w4b:Nat,
                      1,1,nilAM,rb("door","unlock"),5,0,
                      2,2,nilAM,rb("door","lock"),5,0,
                      redirect("dev1","dev2")       --- change tgt
                      unredirect("dev1","dev2")) .  --- change src
Figure 7mod successful attack
We search for a final state in which dev0 has a response from dev1 that the door is locked while in fact the door of dev1 is unlocked, dev1 sent no response, and dev2 did send a response.
              search {caFig7mod(5,0)} =>!
              {c:Conf ["eve" | a:Attrs caps(mtC)]} such that
              not(hasRspTSnt(c:Conf,"dev1","dev0","getN0")) and
              hasRspTSnt(c:Conf,"dev2","dev0","getN0") and
              checkRsrc(c:Conf,"dev1","door","unlock") and
              hasGetRsp(c:Conf,"dev0","dev1","getN0","lock") .
There are 4 solutions out of 33 states visited. In the dialected case there are no solutions among 21 states visited. This is because the GET request is dialected using the secret shared by dev0 and dev1, and dev2 can not decode the message, so it is dropped.
Appendix 0.C Bisimulation proofs
We consider CoAP system configurations sysC of the form
   { [eid-j | devatts-j ] | 0 <= j < k
     net(dmsgs0,dmsgs1)
     ["eve" | kb(dmsgs) caps(caps)]
and dialected CoAP system configurations, dsysC, of the form
   { [eid-j | conf([eid-j | devatts’-j]
                   net(dmsgsO-j,dmsgsI-j))
              ddevatts-j ]
     | 0 <= j < k
     net(ddmsgs0,ddmsgs1)
     [eve | kb(ddmsgs) caps(capsd)]
In Section 7.3 we defined transforms relating initial CoAP system configurations and dialected versions D(sysI) and UD(dsysC). We use these transforms to define a simulation relation $\sim$ and prove two stuttering bisimulation relations:
  • 1.
    $\mathtt{sysI}\approx D(\mathtt{sysI})$ for initial sysI with attack capabilities restricted to drop or delay.
  • 2.
    $\mathtt{dsysI}\approx\mathtt{UDA(dsysI)}$ for initial dsysI with attack capabilites drop, delay, replay, redirect where UDA extends UD to remove replay and redirect capabilities.
Recall the notion of initial configuration is defined in Section 3.4 definition 1.
Definition 3 ($\sim$-bisimulation)
Given a simulation relation $\mathtt{sysC}\sim\mathtt{dsysC}$ between system configurations and dialected system configurations, we say that $\mathtt{iSys}$ is $\sim$-stuttering bisimilar to $\mathtt{iDSys}$ ( $\mathtt{iSys}\approx_{\sim}\mathtt{iDSys}$) iff the following hold:
  • 1.
    $\mathtt{iSys}\sim\mathtt{iDSys}$
  • 2.
    for reachable configurations such that $\mathtt{sysC}\sim\mathtt{dsysC}$
    • 2a.
      if $\mathtt{sysC}\Rightarrow\mathtt{sysC1}$ then there is some $\mathtt{dsysC1}$ with $\mathtt{sysC1}\sim\mathtt{dsysC1}$ such that $\mathtt{dsysC}\Rightarrow^{*}\mathtt{dsysC1}$, and
    • 2b.
      if $\mathtt{dsysC}\Rightarrow\mathtt{dsysC1}$ then there is some $\mathtt{sysC1}$ with $\mathtt{dsysC1}\sim\mathtt{sysC1}$ such that $\mathtt{sysC}\Rightarrow^{*}\mathtt{sysC1}$.
We say $\mathtt{iSys}\approx\mathtt{iDSys}$ if there is a simulation relation $\sim$ such that $\mathtt{iSys}\approx_{\sim}\mathtt{iDSys}$.
In the no-attacker case, intuitively the simulation relation, $\mathtt{sysC}\sim\mathtt{dsysC}$, holds if the two configurations differ only in that in dsysC there are meta-level attributes and some messages are in the global net and are dialected and some are in a local net while there is only one net is sysC.
Definition 4 (UDX and simulation)
The simulation relation $\sim$ is defined by the property
$\mathtt{dsysC}\sim\mathtt{sysC}\equiv(\mathtt{UDX(dsysC)}=\mathtt{sysC})$
where, using the notation above, the function $\mathtt{UDX}$ extends $\mathtt{UD}$ and the result is derived as follows.
  devatts-j = devatts’-j 0 <= j < k
  caps = capsd
[…]
Note that dmsgsI-j has only delayed messages with target eid-j and delay 0, and dmsgsO-j has only delayed messages with source eid-j.
Lemma 1 (Inverse relation)
The transformations D and UD are inverses. For initial (dialected) systems
$\mathtt{UD}(\mathtt{D}(\mathtt{sysI},\mathtt{ids}))=\mathtt{sysI}$
and
$\mathtt{D}(\mathtt{UD}(\mathtt{dsysI}))=\mathtt{dsysI}$
Theorem 0.C.1 tells us that dialecting preserves properties of the underlying protocol, in this case CoAP, expressible in LTL (with out the next operator). The corollary says that allowing drop and/or delay attack capabilities does not change this. That is, dialecting does not protect against these attack capabilities. Theorem 0.C.2 tells us that dialecting protects CoAP messaging against replay an message redirection.
Theorem 0.C.1 (Stuttering bisimulation in the absence of attacks.)
For any initial system configuration sysI with empty attack capabilities, sysI is $\sim$-stuttering bisimilar to D(sysI) written
$\mathtt{sysI}\approx\mathtt{D}(\mathtt{sysI})$
Proof
Let sysI be an initial system configuration with $\mathtt{caps}=\mathtt{mtC}$ and endpoint identifiers eids. Let dsysI be D(sysI). According to the definition above, it suffices to show, $\mathtt{sysI}\approx\mathtt{dsysI}$ using the relation $\sim$ the following.
  • (1)
    $\mathtt{sysI}\sim\mathtt{dsysI}$
  • (2)
    If sysC is reachable from sysI, dsysC is reachable from dsysI, and $\mathtt{sysC}\sim\mathtt{dsysC}$ then
    • (2.i) if $\mathtt{sysC}\Rightarrow\mathtt{sysC1}$ (application of one rewrite rule) then there is some dsysC1 such that $\mathtt{dsysC}\Rightarrow^{<3}\mathtt{dsysC1}$ and $\mathtt{sysC1}\sim\mathtt{dsysC1}$
    • (2.ii) if $\mathtt{dsysC}\Rightarrow\mathtt{dsysC1}$ then there is some sysC1 such that $\mathtt{sysC}\sim\mathtt{dsysC1}$ or $\mathtt{sysC}\Rightarrow\mathtt{sysC1}$ and $\mathtt{sysC1}\sim\mathtt{dsysC1}$.
(1) is by definition of $\sim$ and lemma 1. (2.i) is proved by cases on the rule applied. Using the notation above, assume the rule applies to endpoint eid-j.
  • crl[devsend]: By similarity, eid-j in dsysC has the same attributes and hence the same application message to send. Therefore the same rule applies to the wrapped eid-j with the same attribute updates and new delayed messages added to the local network.
  • crl[rcv]: The corresponding message in dsysC has the same delay and is either in the local net of eid-j or in the global net. In the former case, the rule applies in dsysC with $\mathtt{dsysC}\Rightarrow\mathtt{dsysC1}$. In the latter case, the rule sequence ddevrcv ; rcv applies with $\mathtt{dsysC}\Rightarrow^{2}\mathtt{dsysC2}$. In either case the new state is given by the function rcvMsg(epid,devatts,msg) in dsysC and sysC. Thus $\mathtt{dsysC1}=\mathtt{dsysC2}\sim\mathtt{sysC1}$.
  • crl[ackTimeout]: By $\sim$ this rule is enabled in endpoint eid-j in dsysC. Since the rule outcome depends only devatts-j we have $\mathtt{dsysC}\Rightarrow_{\mathtt{ackTimeout}}\mathtt{dsysC1}\sim\mathtt{sysC1}$.
  • crl[net]: If the corresponding message is in the global net then $\mathtt{dsysC}\Rightarrow_{\mathtt{net}}\mathtt{dsysC}\sim\mathtt{sysC}$. Otherwise the corresponding message is in the local net and $\mathtt{dsysC}\Rightarrow_{\mathtt{ddevsend;net}}\mathtt{dsysC1}\sim\mathtt{sysC1}$.
  • crl[tick]: By definition of mte, if tick is not enabled in dsysC then the local net stub is not empty, and by $\sim$ the only local messages are outgoing (incoming messages have delay 0). Thus
    $\mathtt{dsysC}\Rightarrow_{\mathtt{tick;ddevsend+}}\mathtt{dsysC1}\sim\mathtt{sysC1}$
(2.ii) is also proved by cases on the rule applied.
  • crl[ddevsend]: A message is moved between a local net and the global net, thus $\mathtt{dsysC1}\sim\mathtt{sysC}$ by just revising the network message correspondence.
  • crl[ddevrcv]: A message is moved between the global net and a local net, thus $\mathtt{dsysC1}\sim\mathtt{sysC}$.
  • crl[devsend]: by similarity, the corresponding rule instance is enabled in sysC and the effect depends only on the base endpoint attributes and the message which are the same in sysC and dsysC. Thus $\mathtt{sysC}\Rightarrow_{\mathtt{devsnd}}\mathtt{sysC1}\sim\mathtt{dsysC1}$
  • crl[rcv]: similar to devsend and (2.i)
  • crl[ackTimeout]: similar to (2.1)
  • crl[net]: by $\sim$ the sysC net also has the message to move.
  • crl[tick]: if mte is a non-zero Nat in dsysC then the local stub networks are all empty and the mte for sysC has the same value. Thus $\mathtt{sysC}\Rightarrow_{\mathtt{tick}}\mathtt{sysC1}\sim\mathtt{dsysC1}$ since the function passTime has the same effect in both systems.
Corollary 1
The bisimulation result of Theorem 0.C.1 extends to the case that caps allows delay and/or drop.
Proof
We only need to consider the rule attack. delay/drop attacks do not concern the message content or the deliverability of dialected messages, thus an attack rule application in sysC has a corresponding application in dsysC (possible combined with a ddevsend instance) and vice-versa, with the result being $\sim$ related configurations.
Theorem 0.C.2
Let DC(sysI,caps) be D(sysI) with caps added to the attacker capabilities of the dialected system configuration for any initial system configuration sysI. Then for sysI with attacker capabilities drop and delay and caps containing at most replay,redirect capabilities
$\mathtt{sysI}\approx\mathtt{DC}(\mathtt{sysI},\mathtt{caps})$
Proof
Imagine execution of dialected systems annotates each message with the number of the step that generated/sent it (the devsend rule). Copies of messages generated by the attack rule using the replay capability are annotated by the creation step of the original and the replay attack step that created the copy; and redirected messages are annotated by the creation step and the redirection step. The CoAP rules are unchanged. Define the function UDA to adapt UD so that the annotation is dropped from single annotation messages and for doubly annotated messages, only the original message (sans annotation) is in the network image. Copies are dropped and redirected messages are dropped from the network image. Furthermore, attacker capabilities are reduced to delay,drop.
Now we show $\mathtt{dsysC}\sim\mathtt{UDA}(\mathtt{dsysC})$ is a stuttering bisimulation on configurations reachable from D(sysI,dcaps) where dcaps contains only replay,redirect capabilities and the attack capabilities of sysI are among drop, delay.
1. For the $\mathtt{dsysC}\Rightarrow\mathtt{dsysC1}$ direction let $\mathtt{sysC}=\mathtt{UDA}(\mathtt{dsysC})$ and show $\mathtt{sysC}\Rightarrow^{*}\mathtt{sysC1}=\mathtt{UDA}(\mathtt{dsysC1})$ If the rule used is not the attack rule using replay or redirect capability or delivery of the target message of such an attack, this follows by the proof of Theorem 0.C.1 and its corollary. We consider the rules in remaining cases.
  • [attack(replay(n))] at step k. Let $\texttt{ddm}^{j}$ be the message being replayed ($j<k$) and $\mathtt{ddm1}^{j,k}=delay(\mathtt{ddm},n)^{j.k}$ be the copy. Then $\mathtt{UDA}(\mathtt{dsysC})=>\mathtt{UDA}\mathtt{(dsysC1)}$ by the net rule since $\mathtt{ddm1}^{j,k}$ is not in the UDA image but the attack moves the original from the first to the second network component.
  • [ddrcv] of a replayed message that is delivered to the wrapped endpoint. If this is the original, then sysC can also receive the message. If this is the copy (but the original has not been received, thus the copy was not further delayed) then again sysC can receive the original.
  • [ddrcv] and drop a replayed message. Either the original or copy has already been received and the replayed message will be discarded by UDA thus $\mathtt{sysC}=\mathtt{sysC1}=\mathtt{UDA}(\mathtt{dsysC1})$.
  • [attack(redirect(epid0,epid1))]. Here $\mathtt{sysC}\rightarrow_{\mathtt{drop}}\mathtt{sysC1}=\mathtt{UDA}(\mathtt{dsysC1})$ since redirected messages are dropped.
  • [ddrcv] of a redirected message. This has the effect of a drop. The redirected message has already been dropped in sysC, and is not in the UDA image. Thus $\mathtt{sysC}=\mathtt{UDA}(\mathtt{dsysC})=\mathtt{UDA}(\mathtt{dsysC1})=\mathtt{sysC1}$.
2. For the $\mathtt{sysC}\Rightarrow\mathtt{sysC1}$ direction, for any dsysC such that $\mathtt{UDA(dsysC)}=\mathtt{sysC}$ the argument that $\mathtt{dsysC}\Rightarrow^{<3}\mathtt{dsysC1}$ with $\mathtt{UDA}(\mathtt{dsysC1})=\mathtt{sysC1}$ is the same as the proof of Theorem 0.C.1 as the same rules are available for sysC.
Corollary 2
With notation as for Theorem 0.C.2 and letting caps be a reactive attacker capability set we have: $\mathtt{sysI}\approx\mathtt{DC}(\mathtt{sysI},caps)$
Proof
Reactive caps have the form mc(tpat,spat,false,acts) where acts is a set of elements of the form act(tp,sp,n) with tpat,tp being message target patterns and spat,sp being message source patterns. An attack using capability mc(tpat,spat,false,acts) at step $k$ with matching target $M^{j}$ generated at step $j<k$ is a set of messages $M^{j},M_{1}^{j,k}\ldots M_{l}^{j,k}$ ($M_{i}^{j,k}$ is the result of action $i$ and behave the same as a delayed or redirected message). Thus the bisimulation argument for reactive attacks is the same as that used to prove Theorem 0.C.2.